Linux Commercial HOWTO

Martin Michlmayr, tbm@cyrius.com, http://www.cyrius.com/tbm v4.78, 14 April 1997 This document contains a listing of commercial software and applications which are offered for Linux. It is maintained by Martin Michlmayr <tbm@cyrius.com>.

Contents

1 About this document 4

2 Copyright Information 6

3 Disclaimer 6

4 Related Information 7

5 Linux for the Commercial Market? 7 5.1 Turnkey systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5.2 Large end-user customers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.3 Specialized markets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5.4 Biography . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

6 Website Development 12 6.1 ASWedit, HTML editor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 6.2 Empress DataWEB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 6.3 LinkScan 2.0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 6.4 TalentSoft Web+ (WebPlus) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 6.5 VirtuFlex 1.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 6.6 ThreadTrack and WebTailor from Webthreads. . . . . . . . . . . . . . . . . . . . . . . . . . . 20

7 Databases 21 7.1 c-tree Plus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 7.2 Empress . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 7.3 Essentia . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 7.4 FairCom Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 7.5 Just Logic/SQL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 7.6 KE Texpress . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 7.7 Qddb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 7.8 Raima Database Manager++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 7.9 Empress Embedded RDBMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 7.10 SOLID Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 7.11 Velocis Database Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 7.12 Yard SQL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

8 Data Visualisation and CAD 40 8.1 IDL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 8.2 Megahedron . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 8.3 Tecplot 7.0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 8.4 VARKON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 8.5 XVScan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

9 Development Tools 47 9.1 Amzi! Prolog & Logic Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 9.2 Basmark QuickBASIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 9.3 Dynace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 9.4 Absoft FORTRAN 77 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 9.5 Finesse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 9.6 ISE Eiffel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 9.7 EiffelBench . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 9.8 IdeaFix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 9.9 Khoros Pro 2.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 9.10 MetaCard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 9.11 ObjectManual Release 3.0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 9.12 r-tree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 9.13 sdoc (Source Documenter) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 9.14 SEDIT, S/REXX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 9.15 ST/X (Smalltalk/X) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 9.16 tdb (Tcl Debugger) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 9.17 tprof (Tcl Profiler) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 9.18 View Designer/X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 9.19 XBasic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

10 Emulation 75 10.1 Emulus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 10.2 Executor 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 10.3 Wabi 2.2 for OpenLinux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

11 Financial Software 79 11.1 BB Stock Pro and BB Stock Tool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 11.2 TimeClock v1.07 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

12 Libraries 81 12.1 INTERACTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 12.2 Matrix<LIB> - C++ Math Matrix Library . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 12.3 PKWARE Data Compression Library for Linux . . . . . . . . . . . . . . . . . . . . . . . . . . 84 12.4 SIMLIB IG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

13 Mathematics 87 13.1 Maple V Release 4 - The Power Edition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 13.2 MATCOM and MATCOM MATH LIBRARY . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 13.3 MATLAB and Simulink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

14 Network Servers 92 14.1 Aventail Internet Policy Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 14.2 Critical Angle X.500 Enabler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 14.3 DNEWS News Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

15 Office Tools 95 15.1 The American Heritage Dictionary Deluxe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 15.2 Applixware Office Suite for Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 15.3 D.M.S. Document Management System . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 15.4 HotWire EasyFAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 15.5 NExS, the Network Extensible Spreadsheet . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 16 Text Processing 102

16.1 Edith Pro for X11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

17 System Administration Tools 104 17.1 PerfectBACKUP+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 17.2 Venus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

18 X Windows Related Products 106 18.1 AcceleratedX Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 18.2 BXwidgets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 18.3 BXwidgets/DB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 18.4 AcceleratedX CDE Business Desktop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 18.5 AcceleratedX OSF-Certified Motif . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 18.6 AcceleratedX SGI-Certified OpenGL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112

19 Other Software 114 19.1 Clustor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 19.2 FootPrints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 19.3 Aladdin Ghostscript . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 19.4 LjetMgr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118

20 Free Software for Commercial Hardware 119 20.1 Stallion Technologies Multiport Serial Boards . . . . . . . . . . . . . . . . . . . . . . . . . . . 120

1 About this document

This is the Linux Commercial HOWTO. It contains a listing of commercial software which is available for Linux. The Linux Commercial HOWTO doesn’t contain any information on Linux distributions – this is covered by the Distribution HOWTO. This document was originated by Harald Milz <Harald.Milz@Linux.org>. It is now maintained by Martin Michlmayr <tbm@cyrius.com>. If you need to know more about the Linux Documentation Project or about Linux HOWTO’s, feel free to contact the supervisor Greg Hankins <gregh@sunsite.unc.edu>. Greg Hankins will post the listing to several national and international newsgroups on a monthly basis. In addition, the Linux Commercial HOWTO can be found on the World Wide Web at http://www.cyrius.com/tbm/Commercial-HOWTO. New versions of the Linux Commercial HOWTO are always placed at this site first, so please be sure to check if the copy you are reading is still up to date! The Linux Commercial HOWTO is not a forum for product announcements or marketing hype; it is a service for potential customers and the whole Linux community. Resellers will not be listed; the list is for companies who produce their software themselves. Two main goals are being aimed at:

• It shall help companies who want to run Linux to find software solutions and applications. The international distribution of this list will enhance the contact opportunity.

• It is meant to prove the commercial useability of Linux and thus to encourage other vendors to port

their software as well. Companies and developers who are offering their products for Linux and interested in joining the Linux Commercial HOWTO are invited to fill out the following form and contact me at tbm@cyrius.com. This HOWTO contains tabular entries for each product (example follows). The entry format is similar to the Linux Software Map (LSM) entry (field/stanza lengths are arbitrary). If you want me to add your entry please keep short, otherwise I’ll have to shorten your data. Furthermore, please send me plain ASCII data; no HTML, and no PostScript.

     Category:
             Databases, Data Visualisation, Development tools, Financial
             Software, Mathematics, Network Management, Text Processing,
             X Windows or Other Software.
     Name:
     The name of your application.
     Description:
     Short description of the package, just the basic functionality.
     Distribution media:
     Licensing policy:
     Whatever applies.  Is there a free demo or shareware version available
     via FTP or WWW?  Where?
     OS provisions:
     Kernel version, XFree86 version, Motif version, RAM, harddisk usage, etc.
     Documentation:

     Printed documentation, page number, online help, language.
     Extra features and add-Ons:
     (and their prices)
     Price range:
     Whatever it costs.
     Vendor:
     Address:
     Phone: (U.S. and Canada: if you like to be reachable internationally,
            please don't enter only a +1 800 or +1 900 number)
     Fax:
     EMail:
     URL:
     Contact:

2 Copyright Information

This HOWTO is Copyright \begin{matrix}c \\ ◯ \end{matrix} 1996, 1997 by Martin Michlmayr.

A verbatim copy may be reproduced or distributed in any medium physical or electronic without permission of the author. Translations are similarly permitted without express permission if it includes a notice on who translated it.

Short quotes may be used without prior consent by the author. Derivative work and partial distributions of the Linux Commercial HOWTO must be accompanied with either a verbatim copy of this file or a pointer to the verbatim copy.

Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions.

In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs.

We further want that all information provided in the HOWTOs is disseminated. If you have questions, please contact Greg Hankins, the Linux HOWTO coordinator, at gregh@sunsite.unc.edu.

3 Disclaimer

This HOWTO is not actually a HOWTO in the sense of the Linux Documentation Project. Instead, it is an instrument to investigate the commercial Linux opportunity and to list applications which were already ported and marketed in a native Linux version. As a software vendor, you probably know that you can alternatively offer Linux users a statically linked SCO version of your application which would probably run under the iBCS2 emulator (albeit with a small performance penalty and higher memory requirements). Such applications will not be listed here.

I will not select nor deselect any particular product. Instead, everyone who wants to have her product included will be serviced. However, I reserve the right to shorten individual entries to keep things in shape.

If you don’t find a particular product or vendor in this list, this is probably due to one of the following reasons:

• I never heard of that product or vendor and thus didn’t try to get in contact.

• I did get in contact, but the vendor didn’t answer yet.

• I did get in contact, but the vendor stated positively that he doesn’t sell his product for Linux (yet).

In any case, please get in contact if you feel someone’s missing; also if you discover any errors in the file.

Sometimes two vendor’s addresses are mentioned in the “Vendor:” field. In these cases I received the information from the German subsidiary/distributor. The original manufacturer’s address is always mentioned first.

4 Related Information

There’s another document which covers commercial Linux software. It is maintained by Alan Cox <Alan.Cox@Linux.org> and can be found at http://www.uk.linux.org/LxCommercial.html.

In addition, feel free to visit the LinuxMall where you can order most products presented in this HOWTO.

5 Linux for the Commercial Market?

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1996 iX Multiuser Multitasking Magazin Courtesy of iX Multiuser Multitasking Magazin!

People keep asking me “When is Linux going to be ready for the commercial market”. I guess the first thing to discuss is what is meant by ”commercial” in this context.

Some CD-ROM vendors have put the word ”commercial” in their name, only to have the technical people think their product is good only for use by banks and insurance companies. Other people look at their product with disdain and say that “Linux is not commercial quality”, because it is missing some feature they need, or they feel it is unstable (usually without ever trying it even one time).

To me, the word “commercial” has lost as much meaning in the marketplace as some of the other buzzwords:

• “Open” vs “Proprietary” (People are now saying UNIX is “Proprietary”)

• “Workstation” vs “PC” (What is a workstation, anyway?)

In the old days of computing the commercial market was banks, insurance companies and business-oriented facilities where the use of COBOL or RPG was the mainstream. The technical market was scientific, engineering and manufacturing where FORTRAN and assembly language was used. Somewhere along the way the term “commercial” seems to have gotten twisted around to mean “ready for the mass market”, versus “ready for hackers”.

For the purpose of this article I will take the second meaning, and address whether I think Linux is ready for commercial purposes rather than the hobbyist and hacker market, and ready for the mass market rather than limited markets.

For those of you who hate reading long articles, or who are short on time, let me give you my conclusion right now. Then you can go out and drink beer or other fun activities:

“Yes, Linux is ready for the commercial market...in some cases”.

In order for an operating system to be ready for the mass market it must have several attributes:

• have lots of applications

• be relatively easy to install

• have lots of applications

• be relatively easy to maintain

• have lots of applications

• be relatively easy to use

• have lots of applications

• not crash (much)

• have lots of applications

• be economical

• have lots of applications

But you can eliminate all of these considerations in today’s mass market if only one thing is true:

You have lots of applications.

after all, there would not be 170,000,000 DOS systems in the world if any of the others had to be true.

I almost added that is has to be economical, but history has actually proven me wrong on that. If people added up the total cost of ownership, then Apple would certainly have won over the PC. But people ignore the human costs of someone else (or even themselves) beating their head against the wall trying to get something to work, or the system crashing repeatedly, or the fact that the one keystroke they can hit the easiest (through practice) is

             <CTRL><ALT><DEL>

In the old days people were content to spend several hundreds of dollars on a simple ASCII text editor, or deal with a simple spreadsheet. And it took an act of mangement to get them, with lots of Purchase Orders. Today, they want multi-media integrated with their operating system, and have all the applications available that their neighbor (or boss, or compatriot) has available on their system. And they want to get these applications easily, certainly no harder than to call up on the phone to order them through a catalog, or go down to their corner store to get them.

Now what causes this plethora of applications for an operating system? Ease of programming? Good software development tools? Features inside the operating system? Stability of the interfaces over time?

The answer is “none of these”. While all these attributes may help convince an application developer to port, the one overriding issue is volume of the operating system platform. Again, if MSDOS were compared to MacOS, or even to UNIX and volume were not taken into account, we know which two operating systems would have the most applications, and they would not be from Microsoft.

While it is true that several Linux vendors are working on getting these applications for the mass market (read this ” your mother and father”), the number of applications that run on Microsoft platforms have been estimated as high as 35,000. SunOS has an estimated 10,000 applications, with other ‘commercial UNIX” systems (including Solaris 2.x) much lower in number. It will take the Linux vendors a long time to get the number of applications necessary to hit the really large mass market, particularly if they did not depend on iBCS2 and DOS/Windows compatibility (which could supply a fair number of current applications), but depended on “native” Linux applications.

So applications are king (and queen) for the mass market, and installed base (volume) or the promise of explosive growth (volume) is the key to these. But is the mass market the only “commercial” market? The answer is “no”. The mass market is a subset (albeit very large one) of the commercial market. So let’s look at what the rest of the commercial market needs. We will look at this by segmenting the market into:

• turnkey systems

• large end-user customers

• specialized markets

5.1 Turnkey systems

When I speak of turnkey systems I typically mean a computer system that has one specific (or not so specific) application that runs on it. Examples of turnkey systems are point-of-sale terminals, reservation systems, CAD systems, etc. But in a larger sense, other applications such as Web servers, nameservers (such as BIND), etc. could also be considered “turnkey”, since they have only a few necessary programs that have to run on the system.

Usually turnkey systems are ones that an Independent Software Vendor (ISV) or Value Added Reseller (VAR) will chose a hardware system, an operating system, port an application to it, then duplicate that system 500 to 1000 times without change to the basic application. These ISVs and VARs will try to chose the lowest cost solution to fit their customer’s needs.

Linux is perfect for these types of applications. The operating system is stable enough for the developer to port their application and test the application fully. Once it is fully tested and stable, the entire package is “frozen” and duplicated any number of times for the end customer.

Since the operating system may be freely copied, and it runs on inexpensive hardware, their variable costs are minimal. Even a developer who is not familiar with the Linux system (so they need help getting it running on a platform) will quickly pay back the porting and system programmer costs they accrue by not paying $200-$500. per license for the operating system. Plus they have all the source code for the entire system, in case they run into trouble later on. You can buy a lot of Linux support for $200-$500K.

As I said before, I include Internet Service Providers (ISPs) as part of this “turnkey” environment, for both external internet and internal internet. Why overload your expensive, high-powered, highly complex general-purpose server to do Web serving when a smaller, simpler box can offload it? Why not run your NIS slaves on a Linux box? Or perhaps your BIND server?

In the early days of Digital UNIX (known then as DEC OSF/1) we did not have very many applications. In fact we had none. The marketing staff came to me with sad faces asking if it was possible to sell an operating system that had no applications. I invented a term called “Turbocharging”, which allowed a Digital UNIX system using the speed and power of the Alpha processor (as well as the throughput of our networking devices) to offload NIS, NFS, BIND and other services from people’s overloaded, slower SPARC machines. We also showed people how they could use the rsh(1) command to allow the Alpha to do a portion of their very CPU intensive processing while delivering the result back to the SPARCs on their desk. This allowed the SPARCs to work more on applications and less on the other “system administration” tasks that they were performing. We sold lots of Digital UNIX systems based solely on executing those tasks. Today, of course, Digital UNIX has a lot more applications, and particularly very large memory databases that are extremely fast. But the same principle applies. The database engine runs on the Alpha system, supplying data to the slower SPARC engines as a “Turbocharger”. I could see Linux systems headed in the same direction, following the same path.

5.2 Large end-user customers

Very large customers often have their own home-grown applications which they need to deploy across a wide network of people. Or they can have management dictate a certain suite of applications, which then can be ported to Linux. Since these customers are so large, their operating system costs are huge, and utilizing the savings using the Linux operating system they may completely cover the expenses of porting their software.

Or these very large customers may “influence” their layered product providers to port to the Linux platform. Finally, they may even change some of their computing habits (to use existing programs) if the cost savings are enough to warrant it.

Companies like Caldera are creating a suite of applications and approaching these very large customers to show them the operating system savings that they can achieve if they switch to Linux. While it is true that every application the customer could ever conceive of running may not run on Linux, by using the native applications, the iBCS2 applications, the DOSEMU applications, and applications that run under WABI, a nice suite of applications could be built to solve their needs.

5.3 Specialized markets

Finally there are what I call “specialized markets”. Markets that might buy Linux simply because it is Linux, and not because of the application suites that it provides.

In the education field there are three main markets:

• administrative

• “campus computing”

• computer science education

The administrative part is the “business” aspect of the market. They are looking for easy-to-use systems that can also handle complex administrative tasks that might cover a community the size of a small city.

The “campus computing” is the supply of computing power and service for majors of all types, web services and research into non-computer science (for example, molecular modeling) research.

Finally there is computer science education, both on the undergraduate and graduate level, as well as research into computer science.

While the administrative sub-market typically relies more on shrink-wrapped applications, the other two rely on them to a lesser extent (with the computer science education market relying the least). The other two markets can utilize a lot more of the freeware and shareware applications that are already ported to Linux. This gives them a very low-cost (from a software perspective) platform while allowing them to see and (often) modify the source code for the applications they use.

More importantly, in the computer science research area, the results of the research can be freely distributed to others working in the field, or even published as source code to illustrate the results. This can not be done with “commercial” operating systems.

Some universities are utilizing Linux more and more to run their campus. From a “commercial” standpoint, their needs are the same as many large businesses. Students graduating from college will know about Linux, and bring the word to their future employers.

Finally, there is the computer hobbyist and software developer market. I relate this market to the amateur radio market. In the amateur radio market the radio is often used to simply talk to other people, but at the same time the users investigate new ways of using radio, and improving it. Many electrical engineers started out as amateur radio users. So it can be with Linux, since for the first time both the prices of the hardware and the prices of the operating system source code are within the reach of mortal people.

In conclusion, I feel that Linux does have the items needed for several types of “commercial” uses:

• stability and quality

• low variable costs for turnkey applications

• explosive growth in volume to attract ISVs What Linux really needs is for the “commercial” community to understand what is going on, and to embrace it where it will be useful. This will increase the volume numbers even more, which will attract more applications.

Along these lines I would like to “advertise” a joint effort of USENIX and Linux International to happen in January of 1997 in Anaheim, California of the United States. There will be a joint USENIX/Linux development conference, and while a certain part of the Linux conference will be oriented towards the development of the Linux operating system, the bulk of the conference will be oriented towards application developers and marketing people, to better understand the Linux operating system and how to sell their applications and services into the Linux market. We hope to show ISVs, VARs, resellers and distributors how they can make money by selling their applications and services on top of the Linux operating system.

To join in this effort, please view http://www.usenix.org/linux.html and encourage your VARs, ISVs and resellers (or even yourself!!) to attend. We call the conference ”USELINUX”, and we hope that we can prove to people at that event that it is time to just USE LINUX.

5.4 Biography

Jon “maddog” Hall is a Senior Leader in the Digital Equipment Corporation UNIX group. He has been in the computer industry for twenty-five years, UNIX for sixteen years and has guided the emergence of six operating systems, including Alpha Linux. He has an MS in Computer Science.

6 Website Development

6.1 ASWedit, HTML editor

Description:

ASWedit is a commercial, comprehensive and easy to use HTML and text editor for X Window System and Motif. It offers three independent modes: a plain text editing mode and two context-sensitive, validating modes for authoring of HTML documents as used on the Internet and Intranets. The two HTML modes are: standard and experimental.

Distribution media:

3.5” diskettes, 4-mm DAT, 1/4” and 8mm tapes, Internet (FTP).

Licensing policy:

Per machine basis. The number of users that can run the software on the licensed computer is unlimited. A version of the program, called asWedit, is available for free for students and staff in education and charitable non-profit organizations, and for free evaluation by individuals and commercial organizations. It is available via FTP from many archives. See http://www.advasoft.com/asWedit.html for details.

OS provisions:

Linux 1.2.13 or higher (ELF), X11R6, Motif 2.0 (not required if the statically linked version is used), 5 MB of RAM, 1.5-3.5 MB hard disk usage.

Documentation:

Printed documentation, online help, language: User’s Guide (44 pages), HTML 3.2 extended, Reference Manual (89 pages), HTML 3.2 experimental, Reference Manual (106 pages). Online, context sensitive, hypertext help - 560 KB. Localized resource files are available for the following languages: English (default), Czech, Danish, Dutch, French, German, Polish, Portuguese, Spanish and Swedish. Commands and messages are localized for each language but the online help and documentation are only available in English.

Product support:

The license includes free product upgrades by FTP for a period of one year. Extra features and add-ons:

Can work with external Unix filters. Highly customizable. Supports four different browsers for previewing.

Available since:

July, 1995. Countries with distribution:

World wide. Price range:

US$149. Quantity discounts are available. Vendor:

AdvaSoft Ltd.

     30 Hatch Road
     London SW16 4PN
     England

Phone:

+44 181 251 0033 Fax:

+44 181 251 0011 EMail:

as@advasoft.com URL:

http://www.advasoft.com Contact:

Andre Stochniol Last modified:

March 28, 1997.

6.2 Empress DataWEB

Description:

Empress DataWEB allows users to easily and rapidly build dynamic, interactive, database-fed web applications. No special programming language needs to be learned; developers of applications simply can use HTML with the Empress extensions for accessing the RDBMS.

Distribution media:

CD-ROM. Licensing policy:

Please contact vendor for evaluation copies. OS provisions:

16 MB of RAM; 60 MB harddisk space (additional requirements: web browser required). Documentation:

Online help. Product support:

Full technical support available, priced separately. Extra features and add-ons:

Extra features such as an HTTP server and other tools to facilitate the creation of web applications come with the software package.

Available since:

April, 1996. Countries with distribution:

World wide. Price range:

Please contact vendor. Vendor:

Empress Software Inc.

     6401 Golden Triangle Drive
     Greenbelt, MD 20770
     U.S.A.
     3100 Steeles Avenue East
     Markham, ON L3R 8T3
     Canada

Phone:

+1 301 220 1919 (USA), +1 905 513 8888 (Canada) Fax:

+1 301 220 1919 (USA), +1 905 513 1668 (Canada) EMail:

sales@empress.com URL:

http://www.empress.com Contact:

Dick Naedel Last modified:

March 25, 1997.

6.3 LinkScan 2.0

Description:

LinkScan operates on Unix an NT servers on both Internets and Intranets, LinkScan can test over 40,000 links per hour because it is the only link checker that uses multi-threaded simultaneous processing. LinkScan has been tested on web sites with over 45,000 pages and more than 80,000 links. LinkScan also produces two revolutionary types of maps of web sites. LinkScan’s SiteMap enables the user to produce a site map that includes every link on a web site arranged in a hierarchical format that resembles a book’s table of contents. LinkScan’s TapMap is an expandable and collapsible site map that allows viewers to tap down through the various and multiple levels of a web site to quickly and easily navigate and explore the web site by tapping on a few control icons.

Distribution media:

Internet (WWW). Licensing policy:

A license is required for each server on which the product is used. Free evaluation copies of LinkScan may be downloaded from our web site.

OS provisions:

Requires Perl 5 or higher. Documentation:

Complete documentation and a comprehensive FAQ may be read at our web site and/or downloaded. Product support:

This product is continuously updated and maintained. Prompt responses to all inquiries and problems via EMail or telephone as required. No fees for support.

Extra features and add-ons:

Within a web site, LinkScan offers the following features:

• Scans for missing HTML documents, images and other files

• Validates all internal hyperlinks

• Checks all name tags and references

• Creates two types of site maps or table of contents suitable for publication

• Discovers orphaned files

More importantly, LinkScan will probe each one of your external hyperlinks to other web sites. Scanning these links on a regular basis is essential because you do not have any control over those remote sites. LinkScan detects many different errors most of which will appear as ”broken links” and frustrate your visitors. These conditions include:

• Servers that are not found (no DNS entry)

• Documents that are not found

• URL’s that have relocated

• URL’s that cannot be accessed (e.g. unauthorized or forbidden)

• URL’s that time out

Available since:

January 7, 1997. Countries with distribution:

World wide. Price range:

Licenses for charitable non-profit organizations; or educational institutions (K-12, junior college, college, or library); or individuals for personal use are US$49.95 each. Licenses for government agencies; or commercial enterprises with more than three employees (including officers) or over US$200,000 a year in gross income or are incorporated are $495 each.

Vendor:

Electronic Software Publishing Corporation

     1504 \#8-00200 Main Street
     Gardnerville, NV 89410-5273
     U.S.A.

EMail:

ken@elsop.com URL:

http://www.elsop.com Contact:

Ken Churilla Last modified:

March 26, 1997.

6.4 TalentSoft Web+ (WebPlus)

Description:

TalentSoft Web+ is a development tool dedicated to developing web-based client/server applications without writing low level CGI programs. Web+ enables rapid and easy creation of highly functional web pages which integrates with databases, file systems, EMail, Java applets, your legacy applications (EXEs, DLLs), and communicates with other TCP/IP applications using sockets. Web+ works with all popular web severs, databases, and operating systems and integrates closely with Netscape web servers via NSAPI and CGI. Web+ also acts as a multi-threaded web middleware that integrates the web servers with databases, EMail, TCP/IP sockets, and other applications. Please check out the ”Teach Me Web+” link on our web site for the coolest interactive tutorial with hands on exercises. We are proud to be the first to provide dynamic code interpretation on the Internet! Now you may start writing and running your own web application without buying or installing Web+.

Distribution media:

3.5” diskettes, CD-ROM and Internet (FTP and WWW). Licensing policy:

Free evluation copy available at http://www.TalentSoft.com. Trial out version notices displayed by ill not time out.

OS provisions:

16 MB RAM, 5 MB HD. Documentation:

Both Printed documentation (about 200 pages) and HTML online help. English only. Product support:

Free EMail, web+ conference, and phone support. Training available for a fee.

Available since:

February 1, 1997. Countries with distribution:

USA, UK, Hong Kong, China, Singapore. Price range:

US$49 to $495 Vendor:

TalentSoft / Talent Information Management, LLC.

     P.O. Box 2997
     Minneapolis, MN 55402
     U.S.A.

Phone:

+1 612 338 8900 Fax:

+1 612 904 0010 EMail:

info@TalentSoft.com URL:

http://www.TalentSoft.com Contact:

Victor Tong, Robert King, Tony Tong Last modified:

February 8, 1997.

6.5 VirtuFlex 1.1

Description:

VirtuFlex 1.1 is a web application builder for adding dynamic functionality to a web site. VirtuFlex provides the power to transform web sites into live applications by integrating databases, fax, EMail and pagers with the web. VirtuFlex provides sophisticated functionality that can be added to web sites by any HTML developer. VirtuFlex is reusable, modular, easy to use, high performance and provides DB connectivity almost any database. The componenets of VirtuFlex are a macro language, macro processor, DB server and pre-built template packs.

Distribution media:

Internet (WWW).

Licensing policy:

VirtuFLex is licensed on a per domain basis. A free evaluation copy is available on our web site. OS provisions:

Linux ELF binary format (1.2.x kernels or higher). VirtuFlex runs on standard Unix workstations with 8 MB of RAM minimum, 16 MB recommended.

Documentation:

Available for download from our web site, English. Product support:

Basic support four hours. Other support options available - call for details. Extra features and add-ons:

Pre-built Template-Paks come free with VirtuFlex for shopping carts, threaded discussion groups, database application builder, quizzes and surveys, web spiders and banner rotators. Additional Template-Paks are added on a regular basis.

Available since:

1996. Countries with distribution:

World wide through UniDirect, Soft Export and the Internet. Price range:

US$995, educational discounts available. Vendor:

VirtuFlex Software Corp.

     930 Massachusetts Ave.
     Cambridge, MA 02139
     U.S.A.

Phone:

+1 617 497 8006 Fax:

+1 617 492 0486 EMail:

comments@virtuflex.com URL:

http://www.virtuflex.com Contact:

Dan Housman Last modified:

March 17, 1997.

6.6 ThreadTrack and WebTailor from Webthreads.

Description:

ThreadTrack and WebTailor are lightweight browser and server independent CGI script packages, developed under Linux, that add state to web servers. ThreadTrack is used for tracking the activity of individual visitors to a web site, and WebTailor is used to dynamically modify the content of a web site in response to a visitor’s profile or actions. WebTailor uses a simple server-side scripting language to modify the site’s content. The language, targeted to non-technical web designers, is easy to learn and use. For the more technical, the CGI interface has been expanded to enable parameter passing on a per visitor basis between scripts running on different pages. ThreadTrack tags individual visitors to a web site with a unique identifier that remains with them for their visit. Each visitor’s session is recorded click-by-click, so a database of aggregate and individual activity is available for reports. Comprehensive reporting is included, and the data (dBase III) is easily transferrable to custom reporting packages.

Distribution media:

Internet (WWW). Licensing policy:

30 Day fully functional eval available from http://www.webthreads.com OS provisions:

Any version of Linux on x86, a.out and ELF. Documentation:

Web site and online provided with the package. Product support:

EMail support. Extra features and add-ons:

Msql interface. Registration site management addition. Available since:

June, 1996. Countries with distribution:

World wide.

Price range:

ThreadTrack starts at US$295. WebTailor starts at US$895. Vendor:

Webthreads, L.L.C.

     1919 Gallows Road, 10th floor,
     Vienna, VA 22182
     U.S.A.

Phone:

+1 703 848 9027 Fax:

+1 703 848 2444 EMail:

info@webthreads.com URL:

http://www.webthreads.com Contact:

Gavin Sutcliffe Last modified:

March 18, 1997.

7 Databases

7.1 c-tree Plus

Description:

Based on advanced B+tree (balanced) algorithm, c-tree Plus API handles all aspects of database I/O. Program single user or multi-user non-server applications royalty free and migrate existing c-tree Plus applications to the FairCom Server by recompiling.

Distribution media:

3.5” diskettes and CD-ROM. Licensing policy:

c-tree Plus is licensed on a per programmer basis. Single-user, multi-user and multi-user non-server royalty-free distribution. Contact FairCom for possible restrictions: general purpose database and application development systems prohibited.

OS provisions:

3 MB hard drive space; 128 KB of RAM. Documentation:

Printed manuals distributed with product; full online documentation with CD-ROM, available in English and Japanese.

Product support:

Three months of full technical support from purchase date. Unlimited technical support and product updates available thereafter with c-tree Plus maintenance program.

Available since:

October, 1995. Countries with distribution:

World wide. Price range:

US$895. Vendor:

FairCom

     4006 W. Broadway
     Columbia, MO 65203-0100
     U.S.A.

Phone:

+1 573 445 6833 Fax:

+1 573 445 9698 EMail:

Faircom@faircom.com URL:

http://www.faircom.com Contact:

Tamra Brown <tami@faircom.com> Vendor:

FairCom Europe

          Via Patrioti 6
          I-24021 Albino
          Italy

Phone:

+39 35 773 464 Fax:

+39 35 773 806 EMail:

Europe@faircom.com Vendor:

FairCom Japan

          Ikeda Bldg. \#3 4F, 112-5
          Komei-chou, Tsu-city MIE 514
          Japan

Phone:

+81 059 229 7504 Fax:

+81 059 249 723 Vendor:

FairCom do Brasil Ltda. Phone:

+55 14 224 1610 Fax:

+55 14 234 6462 EMail:

Brazil@faircom.com Last modified:

March 14, 1997.

7.2 Empress

Description:

Empress for Linux is a multimedia RDBMS for members of the rapidly growing Linux developer community. Key components of the package include the powerful Empress RDBMS as well as dynamic SQL, Empress 4GL, Empress GUI Builder for rapidly developing graphical front-ends to Empress applications, a WWW HTML toolkit, and a grahical point and click interface to the Empress RDBMS development environment. A streamlined, single-user version of this product, Personal Empress for Linux, is available also.

Distribution media:

3.5” diskettes. Licensing policy:

Free demos are available via FTP. Must contact vendor. OS provisions:

Empress GUI Builder requires OSF Motif version 1.2.4 or 2.0. 16 MB RAM. 80 MB Disk Space. 486 CPU or better.

Documentation:

Printed documentation. Product support:

Full technical support available, priced separately. Extra features and add-ons:

Other features which are components of the package include shared libraries, shared memory, math library functions and a C language interface.

Available since:

December, 1995. Countries with distribution:

World wide. Price range:

Product is priced by number of concurrent users. Please contact vendor. Vendor:

Empress Software Inc.

     6401 Golden Triangle Drive
     Greenbelt, MD 20770
     U.S.A.

     3100 Steeles Avenue East
     Markham, ON L3R 8T3
     Canada

Phone:

+1 301 220 1919 (USA), +1 905 513 8888 (Canada) Fax:

+1 301 220 1919 (USA), +1 905 513 1668 (Canada) EMail:

sales@empress.com URL:

http://www.empress.com Contact:

Dick Naedel Last modified:

March 25, 1997.

7.3 Essentia

Description:

Database Engine. Some Features: Remote databases, client/server, automatic consistency check, incremental backup, mirroring, shadowing, distributable database, journaling, versions, RISE, object oriented DBMC, implements relational model 3 tier client/server architecture, cooperative servers, language independent user-configurable (English, Spanish, Portuguese).

Distribution media:

4/8 mm. DAT, 150/525 MB tape. Academic version available from:

• ftp://ftp.inter-soft.com

• http://www.inter-soft.com/ftp/pub/essentia

Licensing policy:

Commercial. Free version for Linux available. OS provisions:

10 MB disk space. Documentation:

Available in PostScript and HTML.

Product support:

Contact essentia-info@inter-soft.com for more information. Extra features and add-ons:

SQL Server, ODBC Interface for Windows, JDBC Interface, User servers. Available since:

1993. Countries with distribution:

Argentina, Brazil, Russia, Mexico, Venezuela and USA. Price range:

Linux version for free (with some restrictions). Other prices available on request. Vendor:

InterSoft Argentina S.A.

     C\'{o}rdoba 883 9th. Floor
     Capital Federal (1054)
     Argentina

Phone:

+54 1 318 8900 Fax:

+54 1 318 8997 EMail:

info@inter-soft.com URL:

http://www.inter-soft.com Last modified:

April 8, 1997.

7.4 FairCom Server

Description:

High-performance, multi-threaded, transaction processing server. Features include: industrial quality transaction processing, including full commit and rollback; intermediate save points and complete logging; automatic log management; restart/disaster recovery; user passwords; access security and on-line administration; deadlock detection/resolution; read/write locks at the record/individual key level; more.

Distribution media:

3.5” diskettes and CD-ROM. Licensing policy:

The FairCom Server is licensed on a per machine basis. Contact FairCom for specific licensing questions.

OS provisions:

2 MB of RAM. Documentation:

Printed manuals distributed with product; full online documentation with CD-ROM, available in English and Japanese.

Product support:

Three months of full technical support from purchase date. Unlimited technical support and product updates available thereafter with Server maintenance program.

Available since:

October, 1995. Countries with distribution:

World wide. Price range:

US$445-$3745 depending upon platform and number of users. Special licensing and OEM agreements available.

Vendor:

FairCom

     4006 W. Broadway
     Columbia, MO 65203-0100
     U.S.A.

Phone:

+1 573 445 6833 Fax:

+1 573 445 9698 EMail:

Faircom@faircom.com URL:

http://www.faircom.com Contact:

Tamra Brown <tami@faircom.com> Vendor:

FairCom Europe

          Via Patrioti 6
          I-24021 Albino
          Italy

Phone:

+39 35 773 464 Fax:

+39 35 773 806 EMail:

Europe@faircom.com Vendor:

FairCom Japan

          Ikeda Bldg. \#3 4F, 112-5
          Komei-chou, Tsu-city MIE 514
          Japan

Phone:

+81 059 229 7504 Fax:

+81 059 249 723 Vendor:

FairCom do Brasil Ltda. Phone:

+55 14 224 1610 Fax:

+55 14 234 6462 EMail:

Brazil@faircom.com Last modified:

March 14, 1997.

7.5 Just Logic/SQL

Description:

Relational Database Management System Distribution media:

3.5” diskettes. Licensing policy:

Unlimited runtime included. Documentation:

Printed manual, examples. Product support:

EMail, fax or phone. Extra features and add-ons:

• client-server option: lets run applications on several computers running Windows or Linux, that access a central database on a Linux system.

• web-enabling option: passthrough between an HTTP Web server and a Just Logic/SQL database. No C or Perl required. SQL commands are embedded directly within HTML files. Compatible with Netscape, Apache, NCSA and all other CGI-compliant HTTP servers.

Available since:

1993. Countries with distribution:

World wide. Price range:

US$149-$395. Vendor:

Just Logic Technologies

          P.O. Box 63050, 40 Commerce St.
          Nun's Island, Montreal, QC H3E 1V6
          Canada

Phone:

+1 514 761 6887 Fax:

+1 514 642 6480 EMail:

sales@justlogic.com URL:

http://www.justlogic.com Contact:

Luc Vallieres Last modified:

March 16, 1997.

7.6 KE Texpress

Description:

KE Texpress is a high-speed client server database engine that supports object-oriented, relational and free text data structures and associated operations. It is particularly suited to applications with large data sets, complex operations and large numbers of concurrent users.

Distribution media:

Internet (FTP and WWW) and tape. Licensing policy:

Commercial product licensed by number of concurrent users. A 30 day free trial is available at our web site.

OS provisions:

Linux ELF and a.out libraries. Requires about 30 MB disk space. Runs on over 20 other varieties of Unix.

Documentation:

Printed and HTML documentation is available. Product support:

Annual technical support (EMail, fax and phone) and version upgrade contracts. Extra features and add-ons:

• Texhtml WWW module - publish KE Texpress databases on the WWW

• Texql - SQL-like structured query language

• TexAPI - applications programming interface with client libraries for

• Windows (C, C++ and VB), Unix and Macintosh System 7

• TexODBC - ODBC drivers for Windows KE Software has an extensive consulting service assisting clients to develop KE Texpress database applications.

Available since:

1984. Countries with distribution:

USA, Canada, Australia, Hong Kong, Japan and Malaysia. Price range:

US$2,000 to $100,000+. Monthly fee option of $400 for Internet-based database applications which includes server space, domain name, EMail and Internet connection.

Vendor:

KE Software Inc.

     303-601 West Broadway
     Vancouver, BC V5Z 4C2
     Canada

Phone:

+1 604 877 1960 Fax:

+1 604 877 1961 EMail:

info@kesoftware.com URL:

http://www.kesoftware.com Contact:

Andrzej Kowalski Last modified:

February 21, 1997.

7.7 Qddb

Description:

Qddb is fast, powerful and flexible database software that runs on Unix. Some of its features include: Tcl/Tk programming interface, easy to use, you can have a DB application completely up and, running in about 5 minutes, using nxqddb. CGI interface for quick and easy online databases and guestbooks. Fast, and powerful searching capability. Report generator. Barcharts and graphs. Mass mailings with EMail, letters and postcards.

Distribution media:

3.5” diskettes, tape and Internet (FTP and WWW). Licensing policy:

GNU and Commercial versions available. OS provisions:

GNU version comes with source code. Binary packages available for Linux, FreeBSD, and BSD/OS(BSDI) — RPM and buildkit format.

Documentation:

Online documentation and PostScript files available. Printed manuals also available upon request. Product support:

User and programmer support available along with upgrade contracts. Extra features and add-ons:

Tcl/Tk programming interface, CGI interface, report generator, many useful free and commercial applications built with Qddb.

Available since:

1996. Countries with distribution:

World wide. Price range:

See http://www.hsdi.com/orders Vendor:

Herrin Software Development, Inc.

     41 South Highland Avenue
     Prestonsburg, KY 41653
     U.S.A.

Phone:

+1 606 886 8202 Fax:

+1 606 277 3239 EMail:

info@hsdi.com URL:

http://www.hsdi.com Contact:

Eric Herrin Last modified:

April 11, 1997.

7.8 Raima Database Manager++

Description:

Low-level high performance database engine with C API and C++ class library, for embedding in applications. This database is used in thousands of leading commercial applications. Includes source for C++ class library that encapsulates database navigation and object storage and retrieval into C++ classes, adding an object-oriented interface. Supports multiple database models, including relational, network model, and combined.

Distribution media:

3.5” diskettes and tape. Licensing policy:

Pay for development license, distribute runtime copies freely. Documentation:

Extensive documentation available from Raima Corporation, dealing with all aspects of database, C API and C++ class library.

Product support:

Available from Raima Corporation on annual basis, raining available. Extra features and add-ons:

Windows GUI Report Writer. Available since:

1984, originally called db VISTA. Countries with distribution:

Direct in the USA, use distributors internationally - Germany, England, France, Italy, Spain, Netherlands, Russia, Estonia, Argentina, Columbia, Singapore, Malaysia, Australia, Finland, Taiwan, other countries.

Price range:

RDM++ Database Module is US$995 for single user, US$2,395 for multi-user. System is US$1,395 for single user, US$3,395 for multi-user (System includes RDM++ database, QUERY SQL query tool and REVISE database restructuring tool). All versions available with source code for extra cost. Depends on machine class; lowest is US$3,995 without system utilities; US$6,195 with system utilities.

Vendor:

Raima Corporation

     4800 Columbia Center
     701 5th Avenue
     Seattle, WA 98104

Phone:

+1 800 327 2462, +1 206 515 9477 Fax:

+1 206 748 5200 EMail:

sales@raima.com URL:

http://www.raima.com Contact:

Dave Morse Last modified:

March 18, 1997.

7.9 Empress Embedded RDBMS

Description:

The Empress Embedded RDBMS is a toolkit especially designed for developers of complex embedded systems. The power of this toolkit derives from the Embedded RDBMS which can capture and store bulk objects at impressive speeds. The toolkit is also Internet ready with a JDBC/ODBC bridge available.

Distribution media:

3.5” diskettes. Licensing policy:

Please contact vendor for evaluations copies. OS provisions:

32 MB of RAM, 60 MB disk disk space. Documentation:

Printed documentation.

Product support:

Full technical support available, priced separately. Extra features and add-ons:

Extra features included with the toolkit are Empress Report Writer and enhanced Internet capabilities (ability to use Java applets, etc.) via the HTML toolkit.

Available since:

January, 1997. Countries with distribution:

World wide. Price range:

Please contact vendor. Vendor:

Empress Software Inc.

          6401 Golden Triangle Drive
          Greenbelt, MD 20770
          U.S.A.
          3100 Steeles Avenue East
          Markham, ON L3R 8T3
          Canada

Phone:

+1 301 220 1919 (USA), +1 905 513 8888 (Canada) Fax:

+1 301 220 1919 (USA), +1 905 513 1668 (Canada) EMail:

sales@empress.com URL:

http://www.empress.com Contact:

Dick Naedel Last modified:

March 25, 1997.

7.10 SOLID Server

Description:

SOLID Server is a database engine for new applications and products. It is extremely easy to set up, and has a small footprint. SOLID Server is standards-compliant and full of power. It is perfectly suited for distributed use in countless copies. Its maintenance is care-free, requiring minimal or no administrator attention.

Distribution media:

3.5” diskettes and Internet (FTP and WWW). Licensing policy:

Copy-protected. OS provisions:

At least 2 MB RAM, recommended are 8 MB; harddisk about 3-4 MB. Documentation:

English manuals and online help. WWW. Product support:

Available. Available since:

1994. Countries with distribution:

World wide. Price range:

SOLID Desktop US$99, SOLID Server US$199/seat, SOLID Web Engine US$495. Vendor:

Solid Information Technology Ltd

     Huovitie 3
     FIN-00400 Helsinki
     Finland

Phone:

+358 9 477 4730 Fax:

+358 9 477 47 390 EMail:

info@solidtech.com URL:

http://www.solidtech.com Last modified:

March 17, 1997.

7.11 Velocis Database Server

Description:

Velocis Database Server is designed for database application developers who are looking for a high performance client/server or web database engine. Velocis is a scaleable SQL client/server database engine that provides a rich set of architectural choices and APIs including ANSI SQL, SQL C-API, low-level C-API, C++ class libraries, and support for custom APIs. Unlike typical relational client/server database products, Velocis supports both relational and pointer-based network model databases in any combination as well as processing on either side of the client/server equation. The choices of multiple operating platforms, APIs, processing locality (client or server), and database model can be mixed and combined to satisfy the performance requirements of virtually any application.

Distribution media:

3.5” diskettes and tape. Licensing policy:

Pay for development license. Runtimes are requires to distribute applications. Documentation:

Extensive documentation available from Raima Corporation, dealing with all aspects of database, C API and C++ class library.

Product support:

Available from Raima Corporation on annual basis, raining available. Extra features and add-ons:

Windows GUI Report Writer, Raima Object Manager. Available since:

1993, originally called Raima Database Server. Countries with distribution:

Direct in the USA, use distributors internationally - Germany, England, France, Italy, Spain, Netherlands, Russia, Estonia, Argentina, Columbia, Singapore, Malaysia, Australia, Finland, Taiwan, other countries.

Price range:

1-8 users: US$1,995, 1-25 users: US$3,995, unlimited: US$8,995.

Vendor:

Raima Corporation

     4800 Columbia Center
     701 5th Avenue
     Seattle, WA 98104

Phone:

+1 800 327 2462, +1 206 515 9477 Fax:

+1 206 748 5200 EMail:

sales@raima.com URL:

http://www.raima.com Contact:

Dave Morse Last modified:

March 18, 1997.

7.12 Yard SQL

Description:

The YARD company offers the following products:

• YARD-SQL – Relational SQL client/server database with compliance to X/Open XPG4 and ANSI SQL 92 including network support for accessing remote databases

• YARD-ESQLC – Embedded SQL for C

• YARD-ODBC – ODBC interface for MS Windows clients

• YARD-X – Motif client for database access (No development tool)

• YARD-JDBC – Java Interface

Distribution media:

3.5” diskettes and tapes. Licensing policy:

License number and activation key for each product and installation with user dependent licenses. A Private Edition (limited to 1 user and 5 MB database) for non commercial private use and for evaluation is available at ftp://ftp.yard.de.

OS provisions:

Linux 1.2.13 and all following ELF versions; Motif 1.2.3; RAM usage: 1 MB (minimum) for Shared Memory, 500 KB per User; disk usage: YARD-SQL 10 MB, YARD-ESQLC 1 MB, YARD-ODBC 1.5 MB, YARD-NET 1 MB, YARD-X 5 MB.

Documentation:

Printed 600 pages reference and users guide in German. English documentation is available only as PostScript file.

Product support:

Upon request. Extra features and add-ons:

All products also available for other widely distributed Unix systems (e.g. SCO Unix, SPARC Solaris, IBM, HP, SGI).

Available since:

January, 1994. Countries with distribution:

Current world wide distribution through Distributor Ordix Software <info@ordix.de>, Paderborn, Germany.

Price range:

Upon request. For technical matters contact:

YARD Software GmbH

     Wikingerstr. 18
     D-51107 K\"{o}ln
     Germany

Fax:

+49 221 98664 99 EMail:

yard@yard.de URL:

http://www.yard.de FTP:

ftp://ftp.yard.de Contact:

Thomas Schonhoven <thomass@yard.de> For sales matters contact:

ORDIX Software AG

     Westernmauer 12
     D-33098 Paderborn
     Germany

Fax:

+49 5251 1063 99 EMail:

info@ordix.de Last modified:

March 19, 1997.

8 Data Visualisation and CAD

8.1 IDL

Description:

IDL is the pioneering software for data analysis, visualization, and application development. IDL’s features include flexible I/O, 2D plotting, 3D graphics, volume rendering, image processing, mathematics, statistics, a cross-platform GUI toolkit, plus a high-level, array-oriented programming language. Use IDL for visual data analysis, rapid prototyping, or application development. IDL programs are portable across Linux, Windows 3.1, Windows 95, Windows NT, Mac, PowerMac, Unix and VMS. IDL 5.0 will be shipping very soon. To learn more about IDL 5.0 and to request a Pre Release version go to our web site. IDL Student Version Release 5 will be available for Back-to-School 1997. To learn more about the IDL Student Version send EMail to info@rsinc.com.

Distribution media:

CD-ROM and Internet (FTP). Licensing policy:

Free demo CD-ROM available or download via FTP. OS provisions:

IDL supports Linux kernel version 1.2.8 dynamically linked using Slackware 2.3. Statically linked version also included for use with Linux versions less than 1.2.8. 16 MB RAM minimum recommended. 50 MB disk space.

Documentation:

1600+ pages of documentation on paper and hypertext online help. Product support:

Customer support service, maintenance/update service, training and consulting are available. In addition, a Usenet group is devoted to IDL.

Extra features and add-ons:

HDF support, animation, volume slicer, high resolution mapping built-in. Available since:

September, 1995. Countries with distribution:

Ask vendor. Price range:

Contact vendor/distributor. Vendor:

Research Systems, Inc.

          2995 Wilderness Place
          Boulder, CO
          U.S.A.

Phone:

+1 303 786 9900 Fax:

+1 303 786 9909 EMail:

info@rsinc.com URL:

http://www.rsinc.com FTP:

ftp://ftp.rsinc.com Last modified:

March 17, 1997.

8.2 Megahedron

Description:

A flexible and powerful 3D graphics engine controlled by a high-level interpreted language called SMPL. With it, you can learn about 3D, write your own interactive simulations, animations, and ray-tracings. You can even write your own shaders, customize procedural objects and motions, and use up to 64 computers at once for net rendering. Includes executables for Windows 95, Windows NT x86 and Alpha, SGI Irix and Linux.

Distribution media:

CD-ROM, ISO-9660 and Rock Ridge. Licensing policy:

Very flexible. The owner is allowed to run as many copies as they can, on any platforms. Demo version, full documentation and sample images and scripts available on our web site.

OS provisions:

Supports kernel 1.2.8 or 1.3.15. Non-ELF, a.out executable (QMAGIC). 16 MB of RAM minimum recommended. 7 MB free disk space recommended.

Documentation:

In HTML on CDROM. Product support:

Private news server, EMail. Patches on web site. Extra features and add-ons:

Everything included on one CD-ROM. Available since:

August, 1996. Countries with distribution:

World wide, direct and through dealers. Price range:

US$99 Vendor:

Syndesis Corporation

     235 South Main Street
     Jefferson, WI 53549
     U.S.A.

Phone:

+1 414 674 5200 Fax:

+1 414 674 6363 EMail:

syndesis@threedee.com URL:

http://www.threedee.com Contact:

John Foust <jfoust@threedee.com> Last modified:

February 19, 1997.

8.3 Tecplot 7.0

Description:

Tecplot is a plotting program for visualizing and analyzing engineering and scientific data. Tecplot supports XY, 2D, 3D-surface, and 3D volumetric plotting. With Tecplot you don’t have to write a program, just input your data and start visualizing immediately. For more information, see the Amtec Engineering web site.

Distribution media:

CD-ROM and Internet (FTP). Licensing policy:

Either personal licenses (node locked) or network licenses. Evaluation copies available on CD-ROM or by anonymous FTP.

OS provisions:

12 MB minimum, Linux 1.2 or newer (available in both statically linked and dynamically linked versions). Use the statically linked version if your system does not have Motif 1.2 or newer installed. 32 MB RAM is recommended.

Documentation:

Printed User’s and Reference Manuals, full online help. English. Product support:

Update and technical support included with license for first three months. Extra charge (ask for EUSS) after that. Training courses available.

Available since:

V7 since September 1996, V6 since August 1993.

Countries with distribution:

Many, see Amtec web site. Price range:

US$1795 for personal license, US$2395 for single-user network license, US$1795 for each additional user. All prices the USA and Canada only.

Vendor:

Amtec Engineering, Inc

     PO Box 3633
     Bellevue, WA 98009-3633
     U.S.A.

Phone:

+1 206 827 3304, +1 800 676 7568 Fax:

+1 206 827 3989 EMail:

tecplot@amtec.com URL:

http://www.amtec.com Last modified:

February 19, 1997.

8.4 VARKON

Description:

A high level development tool for CAD and Product modeling applications. Interactive parametric modelling in 2D and 3D with object oriented database. High level CAD and modelling language MBS included.

Distribution media:

Internet (WWW). Licensing policy:

Free binary for Linux. OS provisions:

Requires X but not Motif. Needs 3 MB of harddisk for basic installation and 3MB of RAM to run.

Documentation:

400 pages of documentation included in the free version for Linux. Product support:

Service with continous updates and free consultation available for US $75/month. Extra features and add-ons:

Additional plotterdrivers available free of charge. Available since:

July, 1996. Countries with distribution:

World wide. Price range:

Linux version is free of charge. Vendor:

Microform AB

          Henningsholmsgatan 4
          S-703 69 Orebro
          Sweden

Phone:

+46 19314932 Fax:

+46 19314969 EMail:

info@microform.se URL:

http://www.microform.se Contact:

Johan Kjellander Last modified:

February 5, 1997.

8.5 XVScan

Description:

Image scanning and manipulation software for HP ScanJet scanners. Distribution media:

Internet (EMail and FTP). 4mm DDS2 DAT tape or 3.5 diskettes (additional US$15 in the USA, US$25 international).

Licensing policy:

XVScan is based on XV and is not available in demo version due to licensing restrictions. Distributed with full source code.

OS provisions:

Linux, tested with 1.2.x and 2.0.x, requires a recent generic SCSI driver support to be built in (no earlier than 1.1.79). Motif is not required. Any XFree version (X11R5, X11R6).

Documentation:

Online, WWW, English. Product support:

Updates free for first year. Extra features and add-ons:

Also available for HP-UX, BSD/OS 2.2 and 3.0, SunOS 4.1.x, Solaris 2.x, and FreeBSD. Available since:

May, 1995. Countries with distribution:

World wide. Price range:

US$50 for FTP or EMail shipping. Additional US$15 for media in the United States, US$15, internationally. Mastercard, Visa and Discover, and American Express Credit Cards accepted.

Vendor:

tummy.com, ltd.

     3506 Stratton Drive
     Fort Collins, CO 80525-2722
     U.S.A.

Phone:

+1 970 223 8215 Fax:

+1 408 490 2728 EMail:

xvscan@tummy.com URL:

http://www.tummy.com/xvscan Contact:

Sean Reifschneider <jafo@tummy.com> or Evelyn Mitchell <efm@tummy.com> Last modified:

February 19, 1997.

9 Development Tools

9.1 Amzi! Prolog & Logic Server

Description:

Amzi! enables the easy integration of intelligent components with conventional applications allowing you to add logic-bases that give advice, configure and tune systems, diagnose problems, apply business rules, monitor processes and parse documents. Your applications access a logic-base of rules just as a database server accesses records. The rules are expressed in Prolog which has powerful, built-in search and pattern matching capabilities. The Amzi! Logic Server is encapsulated as a C++ Class and C API Interface. You can add your own Prolog functions in C/C++. Includes: compiler, listener, debugger, linker, EXE, generator, call-in/call-out Logic Server API, full documentation, comprehensive Prolog tutorial and lots of sample code.

Distribution media:

Internet. Licensing policy:

The Professional Edition includes an unlimited, royalty-free license. The Personal Edition is limited to distributing applications for non-commercial use only.

OS provisions:

About 350 KB of RAM for typical small application, 3-4 MB disk. Documentation:

HTML format, includes full Prolog tutorial. Product support:

Subscription Plus service provides automatic updates for a full year, $198. Custom development services available.

Linux Support:

No Linux system included. For Amzi! software, free tech support for registered users by phone, fax and EMail.

Available since:

November 3, 1995. Countries with distribution:

World wide. Price range:

Personal (Shareware), $49. Professional, $298. Vendor:

Amzi! inc.

     40 Samuel Prescott Drive
     Stow, MA 01775
     U.S.A.

Phone:

+1 508 897 7332 Fax:

+1 508 897 2784 EMail:

info@amzi.com URL:

http://www.amzi.com Contact:

Mary Kroening Last modified:

March 19, 1997.

9.2 Basmark QuickBASIC

Description:

The Basmark QuickBASIC Compiler is a multi-user IBM-PC BASICA, MBASIC and Microsoft QuickBASIC Compiler designed to provide performance and consistency across a variety of machines (e.g. i386 and i486, Pentium, SPARC, RS/6000, HP PA-RISC) under Unix, AIX, SunOS, Linux, HP-UX and Xenix.

Distribution media:

3.5” diskettes. Licensing policy:

Per machine, no run-time restrictions. OS provisions:

GNU GAS and LD must be installed. Documentation:

400+ page manual, release and installation notes. Product support:

Updates available for US$139 less shipping. Maintenance contracts available. Extra features and add-ons:

C-ISAM (Informix Inc.) Interface in C source code form. Cost is US$35. Available since:

December, 1993. Countries with distribution:

World wide. Price range:

US$695 (less shipping) Vendor:

Basmark Corporation

          P.O. Box 40450
          Cleveland, OH 44140
          U.S.A.

Phone:

+1 216 871 8855 Fax:

+1 216 871 9011 EMail:

jgo@basmark.com (for orders) URL:

http://www.XBasic.com Contact:

Joseph O’Toole (for orders) Last modified:

February 25, 1997.

9.3 Dynace

Description:

Dynace (pronounced like “dynasty” without the “t”) is a preprocessor, include files and a library which extends the C language with advanced object oriented capabilities, automatic garbage collection and multiple threads. Dynace is designed to solve many of the problems associated with C++ while being easier to learn and containing more flexable object oriented facilities. Dynace is able to add facilities previously only available in languages such as Smalltalk and CLOS without all the overhead normally associated with those environments.

Distribution media:

Internet (WWW); including full C source code. Licensing policy:

One license per programmer; Applications are royalty free. Information and demo on http://www.edge.net/algorithms. Free for non-commercial use.

OS provisions:

Any Linux. Documentation:

310 pages; in Postscript, HP PCL, or TeX DVI. Product support:

Dynace-support@edge.net Extra features and add-ons:

Complete Windows development system. Available since:

December, 1993. Price range:

US$599. Vendor:

Algorithms Corporation

          3020 Liberty Hills Drive
          Franklin, TN 37067
          U.S.A.

Phone:

+1 800 566 8991, +1 615 791 1636 Fax:

+1 615 791 7736 EMail:

blake@edge.net URL:

http://www.edge.net/algorithms Contact:

Blake McBride Last modified:

March 15, 1997.

9.4 Absoft FORTRAN 77

Description:

VAX/VMS compatible ANSI FORTRAN 77 compiler and debugger. Distribution media:

3.5” diskettes. Licensing policy:

Single user license; multi-user packages also available. OS provisions:

Linux 1.2.13. Documentation:

300+ pages hard copy documentation. Product support:

No-charge technical support. Available since:

March, 1996.

Countries with distribution:

World wide. Price range:

US$525-$2,000. Vendor:

Absoft Corporation

          2781 Bond Street
          Rochester Hills, MI 48309
          U.S.A.

Phone:

+1 810 853 0050 Fax:

+1 810 853 0108 EMail:

sales@absoft.com URL:

http://www.absoft.com Contact:

Wood Lotz Last modified:

March 25, 1997.

9.5 Finesse

Description:

OSF/Motif GUI for shell scripts. Distribution media:

Internet (FTP). Licensing policy:

Finesse is a fully commercial product for other Unixes where nodelocked licenses are possible. The Linux version is freely available via FTP. Floppy medium is DM 90.

OS provisions:

ELF Libraries. X11R6. No Motif required.

Documentation:

PostScript File. Product support:

No support for free version. Available since:

May, 1995. Countries with distribution:

World wide. Price range:

Free demo. Commercial version with support on request. Vendor:

science + computing GmbH

          Hagellocher Weg 71
          D-72070 T\"{u}bingen
          Germany

Phone:

+49 7071 9457 0 Fax:

+49 7071 9457 27 EMail:

info@science-computing.de URL:

http://www.science-computing.de FTP:

ftp://ftp.science-computing.de/pub/finesse. Contact:

Olaf Flebbe Last modified:

February 25, 1997.

9.6 ISE Eiffel

Description:

ISE Eiffel is a seamless object-oriented development environment. ISE Eiffel provides an integrated solution for software developers through pure object-oriented methods, from analysis and design through code generation, maintenance, and reverse engineering. The components of ISE Eiffel include EiffelBench, EiffelBase, EiffelBuild, EiffelVision, EiffelLex, EiffelParse, EiffelNet, EiffelStore, ObjEdit, EiffelCase, EiffelMath, EiffelWeb, DLE (Dynanic Linking in Eiffel) and SCOOP (Distribution/Concurrency mechanism).

Distribution media:

CD-ROM and Internet (FTP and WWW) Licensing policy:

No run-time fees are required for products, commercial or otherwise, developed with ISE’s technology. In the case of commercial products we simply require the product and its documentation to acknowledge ISE Eiffel clearly. The copyright holder must acknowledge the product’s use/incorporation of ISE Eiffel by (a) Featuring ”ISE Eiffel” prominently in product documentation, and (b) Featuring an ”ISE Eiffel” software window or pop-up message clearly visible during the copyrighted product’s installation process.

OS provisions:

Linux ELF or a.out. Motif version 1.2. Documentation:

Printed documentation. Additional online help. Product support:

Update service, maintenance, training available. Available since:

August, 1994. Countries with distribution:

World wide. Price range:

Starting at US$99. Vendor:

Interactive Software Engineering, Inc.

     270 Storke Road, Suite 7
     Goleta, CA 93117
     U.S.A.

Phone:

+1 805 685 1006 Fax:

+1 805 685 6869 EMail:

info@eiffel.com URL:

http://www.eiffel.com Last modified:

April 8, 1997.

9.7 EiffelBench

Description:

Object-oriented CASE Workbench. Distribution media:

CD-ROM and Internet (FTP and WWW). Licensing policy:

No run-time fees are required for products, commercial or otherwise, developed with ISE’s technology. In the case of commercial products we simply require the product and its documentation to acknowledge ISE Eiffel clearly. The copyright holder must acknowledge the product’s use/incorporation of ISE Eiffel by (a) Featuring ”ISE Eiffel” prominently in product documentation, and (b) Featuring an ”ISE Eiffel” software window or pop-up message clearly visible during the copyrighted product’s installation process.

OS provisions:

Linux ELF or a.out. Motif version 1.2. Documentation:

Eiffel: The Environment. Product support:

Update service, maintenance, training available. Available since:

August, 1994. Countries with distribution:

World wide.

Price range:

Starting at US$99. Vendor:

Interactive Software Engineering, Inc.

     270 Storke Road, Suite 7
     Goleta, CA 93117
     U.S.A.

Phone:

+1 805 685 1006 Fax:

+1 805 685 6869 EMail:

info@eiffel.com URL:

http://www.eiffel.com Last modified:

April 8, 1997.

9.8 IdeaFix

Description:

The InterSoft Development Environment for Applications in Unix, or IdeaFix, is a set of programming tools and utilities designed to provide an integrated environment for programmers as well as end-users. The aim of IdeaFix is to maximize performance and productivity for both. For the end-user, IdeaFix offers a user-friendly interface and an online help system. IdeaFix provides the developer with tools to simplify development of such design criteria as relational databases and modular structured programming.

Distribution media:

4/8 mm. DAT, 150/525 MB tape. Academic version available from:

• ftp://ftp.inter-soft.com/pub/ideafix

• http://www.inter-soft.com/ftp/pub/ideafix

Licensing policy:

Commercial. Free version for Linux available.

OS provisions:

20 MB disk space, and GNU C/C++ compiler. Documentation:

Available in PostScript. Product support:

Contact ideafix-info@inter-soft.com for more information. Extra features and add-ons:

SQL server, Ibuild, Cracker for Windows. Available since:

1986. Countries with distribution:

Argentina, Brazil, Russia, Mexico, Venezuela and USA. Price range:

Linux version for free (with some restrictions). Other prices available on request. Vendor:

InterSoft Argentina S.A.

          C\'{o}rdoba 883 9th. Floor
          Capital Federal (1054)
          Argentina

Phone:

+54 1 318 8900 Fax:

+54 1 318 8997 EMail:

info@inter-soft.com URL:

http://www.inter-soft.com Last modified:

April 8, 1997.

9.9 Khoros Pro 2.1

Description:

Khoros Pro 2.1 is a software development environment with extensive image processing and data visualization capabilities.

Distribution media:

CD-ROM. Licensing policy:

Single user license come with Khoros Pro. For software developers, independent software developer licenses are available. The R&D version of Khoros is available via FTP.

OS provisions:

Linux 2.0.18, gcc 2.7.2, Fortran compiler f77 4.0.1. Documentation:

Printed documentation; Installation Guide and User’s Guide come with the CD-ROM. Four volume Developer’s Manual set available from KRI.

Product support:

Training in Software Development and Digital Image Processing scheduled through 1997. See http://www.khoral.com/training/training.html. Maintenance, technical and engineering support may be purchased from KRI on a negotiated basis.

Available since:

August, 1996. Khoros Pro 2.2 to ship late summer 1997. Countries with distribution:

World wide. Price range:

Single user Khoros Pro 2.1 is US$225. Independent software developer licenses start at US$5,000 and go up depending on platforms and organization structure. Royalty agreements can be negotiated.

Vendor:

Khoral Research Inc.

     6001 Indian School NE Suite 200
     Albuquerque, NM 87110
     U.S.A.

Phone:

+1 505 837 6500 Fax:

+1 505 881 3842 EMail:

info@khoral.com URL:

http://www.khoral.com FTP:

ftp://ftp.khoral.com Contact:

Annie MacFarlane <annie@khoral.com> Last modified:

March 25, 1997.

9.10 MetaCard

Description:

MetaCard is a GUI development and multimedia authoring tool compatible with Apple Corporation’s HyperCard. Anyone can use MetaCard to build GUI applications and hypermedia documents using a powerful, direct manipulation editor and an easy-to-learn scripting language. MetaCard goes beyond HyperCard by including support for color controls and images, vector graphics, scrollbars, and dialog boxes. MetaCard’s scripting language has support for arrays, custom (user defined) object properties, and is based on high-perforance ”virtual compiler” technology. Stacks developed with MetaCard are portable among all popular Unix platforms and Windows 95/NT and can be distributed with without licensing fees or royalties.

Distribution media:

3.5” diskettes and Internet (FTP and EMail). Licensing policy:

MetaCard can be licensed to a single, named individual, but can be used on any machine or combination of machines by that user. Multiple user packages are also available. The save-disabled distribution is available via anonymous FTP from ftp://ftp.metacard.com/MetaCard and ftp://ftp.uu.net/vendor/MetaCard.

OS provisions:

The Linux engine is built on a 1.2.13 ELF system. Minumum 8 MB RAM for runtime, 16MB RAM for development. Minimum 640x480x8 screen resolution for runtime, 800x600x8 for development. Development system requires about 5MB disk space. GUI is Motif-compliant, but does not require Motif libraries.

Documentation:

Complete online documentation is supplied. Printed documentation is an extra-cost option. Product support:

Free EMail technical support, phone support available at extra cost. Extra features and add-ons:

A library version of MetaCard that can be linked directly to C programs (Embedded MetaCard) is available at extra cost.

Available since:

June, 1992. Countries with distribution:

World wide. Price range:

US$995 for single-user all-platform development license. Vendor:

MetaCard Corporation

          4710 Shoup pl.
          Boulder, CO 80303
          U.S.A.

Phone:

+1 303 447 3936 Fax:

+1 303 499 9855 EMail:

info@metacard.com URL:

http://www.metacard.com Last modified:

March 15, 1997.

9.11 ObjectManual Release 3.0

Description:

Automated documentation generator from C++ programs. Generates doucmentation in various formats including HTML, RTF, MIF along with Java classes.

Distribution media:

Internet (FTP and WWW). Licensing policy:

Single, Multiple, Site and Floating Licenses. OS provisions:

Linux ELF. Documentation:

Provided in PostScript along with the distribution. Product support:

Free for 30 days. Then a contract based on six month to a one year cycle. Extra features and add-ons:

New release are released every quarter. Available since:

1995. Countries with distribution:

USA. Price range:

US$3395 for most flavours of Unix. Free of charge for Linux users that work in a non-commercial environment.

Vendor:

ObjectSoftware, Inc.

     3519 Misty Meadow Dr.
     Dallas, TX 75287
     U.S.A.

Phone:

+1 214 373 2021 Fax:

+1 972 662 0756 EMail:

objsoft@netcom.com URL:

http://www.obsoft.com Contact:

Bobby Sardana Last modified:

March 16, 1997.

9.12 r-tree

Description:

r-tree report generator in C source code. Provides complex, multi-line reports by handling virtually every aspect of report generation. The only programming requirement is to call the r-tree report function, which reads c-tree data files, performs calculations, monitors control breaks and accumulators and produces a formatted report. Complete with C source code. Requires c-tree Plus.

Distribution media:

3.5” diskettes and CD-ROM. Licensing policy:

r-tree Plus is licensed on a per programmer basis. Royalty-free distribution. Contact FairCom for possible restrictions.

Documentation:

Printed manuals distributed with product; full online documentation with CD-ROM, available in English and Japanese.

Product support:

Three months of full technical support from purchase date. Unlimited technical support and product updates available thereafter with r-tree maintenance program.

Available since:

October, 1995. Countries with distribution:

World wide. Price range:

US$445.

Vendor:

FairCom

          4006 W. Broadway
          Columbia, MO 65203-0100
          U.S.A.

Phone:

+1 573 445 6833 Fax:

+1 573 445 9698 EMail:

Faircom@faircom.com URL:

http://www.faircom.com Contact:

Tamra Brown <tami@faircom.com> Vendor:

FairCom Europe

          Via Patrioti 6
          I-24021 Albino
          Italy

Phone:

+39 35 773 464 Fax:

+39 35 773 806 EMail:

Europe@faircom.com Vendor:

FairCom Japan

          Ikeda Bldg. \#3 4F, 112-5
          Komei-chou, Tsu-city MIE 514
          Japan

Phone:

+81 059 229 7504 Fax:

+81 059 249 723 Vendor:

FairCom do Brasil Ltda. Phone:

+55 14 224 1610 Fax:

+55 14 234 6462 EMail:

Brazil@faircom.com Last modified:

March 14, 1997.

9.13 sdoc (Source Documenter)

Description:

This program helps to create a complete, indexed documentation from your sources (C, tcl, Perl currently available, other will follow soon). You add documentation information to your scripts using pod (Plain old documentation), the documentation format also used in Perl. The documentation options provided by pod are simple and easy to learn but powerfull enough to create nice and good looking documents. sdoc uses this information to create a pod-document which in turn may be transformed to HTML, LaTeX (and PostScript), nroff or just plain text. The preview function of tdoc lets you create documentation interactively - add or modify your documentation to the source file and check the output for correctness. Here are some highlights:

• Documentation is included in source file, no separate files needed

• Powerfull formatting options combined with easy learning

• Good looking documentation can be created without hassle

• Automated indexing

• Output may be in HTML, PostScript and a lot of other formats (have a look at the perlpod manual page or at available pod converters).

The Tcl Documenter is also available as a commercial version. This version adds the following features:

• Significantly faster • May optionally use Netscape to display preview pages.

• Includes the package concept to further automate file selection.

• Direct connection to emacs for file editing

• Direct creation of postscript output

• Full support, upgrades

• Complete documentation

Distribution media:

3.5” diskettes and Internet (EMail and WWW). Licensing policy:

You can download the public domain version of sdoc from the neosoft tcl archive:

• http://www.neosoft.com/tcl/ftparchive/sorted/development

• ftp://ftp.neosoft.com/pub/tcl/sorted/development

OS provisions:

You need an ELF based system with X11. Documentation:

Context sensitive help and complete documentation in Pod format. Available as PostScript on request.

Product support:

Installation support included, additional support available. Available since:

July, 1996. Countries with distribution:

World wide. Price range:

US$120 for the binary version, US$300 for the source, with a 10% discount for students and educational institutions.

Vendor:

softWorks, Richard Schwaninger

     Theodor-K\"{o}rnerstr. 173
     A-8010 Graz
     Austria

Phone:

+43 316 686590 Fax:

+43 316 686590 EMail:

risc@ping.at URL:

http://members.ping.at/risc Contact:

Richard Schwaninger Last modified:

March 26, 1997.

9.14 SEDIT, S/REXX

Description:

SEDIT is a powerful Unix text editor patterned after IBM’s XEDIT editor. It operates with a GUI under X windows or in character mode from a tty device. S/REXX is a full Unix implementation of IBM’s SAA procedural language except that the numeric digit specification is limited to 15 digits. S/REXX functions as an imbedded macro language for SEDIT as well as providing a powerful modern programming language alternative to shell scripting languages. SEDIT and S/REXX may be purchased in a bundle or individually. See our web site for more information including pricing for Linux, etc. For additional information, please contact one of the local distributors listed on the web site or mail us at sedit@dialup.FranceNet.fr

Distribution media:

4mm, 8mm, 1/4” or 3.5” diskettes and Internet (FTP). Licensing policy:

SEDIT and S/REXX are licensed products. See the above WWW site for details. For demo purposes, a short term license key will be provided on request to anyone who obtains the product via ftp.

OS provisions:

We believe these products will work with any recent stable Linux kernel since 1.1.18 supporting a.out format. Motif is fully bound with SEDIT and SEDIT is distributed with XFree86 3.1 shared libraries for use if your system is not at that level.

Documentation:

Documentation is furnished in PostScript form, in extensive online help files, and in optional printed manuals.

Product support:

Technical support is available for all licensed users. For maintenance, see the above URLs.

Extra features and add-ons:

A REXX interactive graphical debugger is available as an optional feature as are printed manuals, physical media, and on going maintenance. See the above URLs.

Available since:

The original Linux version has been available since 3rd quarter 1995. Countries with distribution:

SEDIT and S/REXX are available worldwide. The URLs should be consulted for current information. Price range:

Pricing information should be requested from the local distributor responsible for your country. An economical 2 user license is available for Linux without media or printed documentation with other options for more users, S/REXX, hardcopy manuals, etc.

Vendor:

Benaroya

     31 Rue de Constantinople
     F-75008 Paris
     France

Phone:

+33 1 47 22 22 13 Fax:

+33 1 47 22 06 17 EMail:

sedit@dialup.FranceNet.fr URL:

http://www.sedit.com Contact:

Robert Benaroya Last modified:

March 19, 1997.

9.15 ST/X (Smalltalk/X)

Description:

ST/X is a complete implementation of the Smalltalk programming language with development environment providing all of the advantages you would expect from Smalltalk: a fully object oriented programming language, a graphical development environment with editors, browsers, symbolic debugger, incremental compiler and just-in-time compilation to machine code as well as an extensive class library.

Distribution media:

CD-ROM.

Licensing policy:

There are non-commercial, evaluation and commercial licences available. There is also demo version on the Internet, although it is quite old, this should be updated at some time in 1997.

OS provisions:

ST/X requires 32 MB or more of memory and 100 MB hard disk space.

Documentation:

The ST/X documentation is available online as hypertext (HTML).

Product support:

eXept Software AG provides mentoring, training, and consulting. Onsite support is also available. For the commercial licence there is an update service and maintenance support.

Extra features and add-ons:

ST/X includes a supplementary stand-alone machine code compiler and the possibility to embed C code directly into Smalltalk methods. This allows the exploitation and use of existing C libraries or a performance improvement in the speed of time critical operations. An integral part of ST/X is an HTML browser which is fully implemented in Smalltalk. It supports network loadable Smalltalk applets and smalltalk scripts embedded in HTML pages. This allows the development of Internet and Intranet applications analogous to Java.

Available since:

ST/X has been available as a beta version on the Internet for quite a while. eXept Software AG has been further developing and marketing ST/X since 1996. Claus Gittinger (the developer from ST/X) is a founding member of eXept.

Countries with distribution:

World wide.

Price range:

There are non-commercial, evaluation and commercial licences available. The commercial licence costs US$3,500. A six month evaluation licence costs US$350. The non-commercial license is available for US$150 plus shipping and handling.

Vendor:

eXept Software AG

  Besigheimerstr. 1
  D-74369 L\"{o}chgau
  Germany

Phone:

+49 7143 870045 Fax:

+49 7143 870048 EMail:

info@exept.de Contact:

David Queeney Last modified:

March 5, 1997.

9.16 tdb (Tcl Debugger)

Description:

This is a debugger for tcl scripts. It allows interactive debugging of tcl/tk programs and speeds up your development cycle. Here are some highlights:

• Step by step execution with breakpoints

• Display and manipulation of local and global variables

• On-the-fly source code modifications and fast programm restarts

• Graphical user interface, easy use

• Easy configuration and installation The commercial version adds the following features:

• Full screen context display

• Hot link to editor (currently emacs)

• Scripting and user extensibility (tcl)

• Plugin support (script, popup)

• A lot faster

• Full support, upgrades

• Complete documentation Distribution media:

3.5” diskettes and Internet (EMail and WWW) Licensing policy:

You can download the public domain version of tdb from the neosoft archive:

• http://www.neosoft.com/tcl/ftparchive/sorted/development

• ftp://ftp.neosoft.com/pub/tcl/sorted/development

OS provisions:

ELF based system with X11. Note that tdb requires changes to the base tcl C-libraries and also depends on dynamic loading. Please check out the info package for further details. You don’ need tcl installed to use the debugger, all necessary code is included. The commercial version may be purchased in source form and support for porting to other platforms is available. Please contact softWorks for details.

Documentation:

Context sensitive help and complete documentation in HTML format. Available as PostScript on request.

Product support:

Installation support included, additional support available. Extra features and add-ons:

Debugger-PlugIns for speacialized debug tasks (e.g. to debug tk widget hierarchies are also available). Available since:

July, 1996. Countries with distribution:

World wide. Price range:

US$450 for a binary version, US$750 for the source code, with a 10% discount for students and educational institutions.

Vendor:

softWorks, Richard Schwaninger

     Theodor-K\"{o}rnerstr. 173
     A-8010 Graz
     Austria

Phone:

+43 316 686590 Fax:

+43 316 686590 EMail:

risc@ping.at URL:

http://members.ping.at/risc Contact:

Richard Schwaninger Last modified:

March 26, 1997.

9.17 tprof (Tcl Profiler)

Description:

This is a profiler for tcl scripts. It allows the creation of profiling data from your tcl programms and later analysis and detailed inspection through a graphical frontend. This tool is especially well suited if you encounter performance problems with your tcl scripts. Instead of rewriting the whole application in C you can isolate time-critical procedures and speed them up (if necessary using C). Here are some highlights:

• Call flow display

• Barcharts for number of calls, cpu time and real time

• Tabular display of profiling data set

• Analyse both tcl procs and commands written in C

• Graphical user interface, easy use

• Simple aquisition of profiling data

• Easy configuration and installation tprof is available both as a public domain and as a commercial version. The commercial version adds the following features:

• Zoom in on interesting subsets of the data

• Printout of data sets

• PostScript (eps) images of charts and tables • Merging of multiple data sets

• Exporting data sets in different formats to inspect them further (eg. sc/xspread, Wingz, DBase, text)

• A lot faster

• Full support, upgrades

• Complete documentation

Distribution media:

3.5” diskettes, Internet (EMail and WWW). Licensing policy:

You can download the public domain version of tprof from the neosoft tcl archive:

• http://www.neosoft.com/tcl/ftparchive/sorted/development

• ftp://ftp.neosoft.com/pub/tcl/sorted/development

OS provisions:

You need an ELF based system with X11. Documentation:

Context sensitive help and complete documentation in HTML format. Available as PostScript on request.

Product support:

Installation support included, additional support available. Available since:

October, 1996. Countries with distribution:

World wide. Price range:

US$250 for the source code version (US$130 for a binary) with a 10% discount for students and educational institutions.

Vendor:

softWorks, Richard Schwaninger

     Theodor-K\"{o}rnerstr. 173
     A-8010 Graz
     Austria

Phone:

+43 316 686590 Fax:

+43 316 686590 EMail:

risc@ping.at URL:

http://members.ping.at/risc Contact:

Richard Schwaninger Last modified:

March 26, 1997.

9.18 View Designer/X

Description:

Motif interface builder that generates C and C++ code. VDX provides WYSIWYG view, widget tree browser, resource editor, widget templates and more tools to design user interfaces.

Distribution media:

Internet (WWW). Licensing policy:

Free demo version. Purchase of license key that avoids restrictions of the demo version. The license key is derived from hardware address of network interface card.

OS provisions:

Linux, ELF, X11R6, Libc.5, 5 MB on harddisk. Documentation:

Online help, English. PostScript manual, German. Product support:

Minor update (1.x to 1.x) free. Free EMail support. Available since:

December, 1996. Countries with distribution:

World wide. Price range:

DM 800 (approximately US$470), Major update (1.x to 2.x) DM 500 (approximately US$295).

Vendor:

Bredex GmbH

     Fallersleber-Tor-Wall 23
     D-38100 Braunschweig
     Germany

Phone:

+49 531 24 33 00 Fax:

+49 531 24 33 099 EMail:

info@bredex.de URL:

http://www.bredex.de Contact:

Regina Heine Last modified:

March 17, 1997.

9.19 XBasic

Description:

XBasic is a comprehensive visual program development environment that integrates a powerful editor, advanced 32/64-bit BASIC compiler, debugger, GuiDesigner and GraphicsDesigner. XBasic is written entirely in XBasic.

Distribution media:

3.5” diskettes and Internet (FTP). Licensing policy:

Each copy is licensed on a per machine basis. OS provisions:

Linux (ELF 1.2.13 or higher). Documentation:

300+ page (8.5 inches by 11 inches) manual. Product support:

US$3 per minute, 15 minute minimum.

Extra features and add-ons:

Enterprise (socketized internet-TCP/IP) edition is available. Available since:

January, 1996. Countries with distribution:

World wide, Price range:

US$795-$955. Vendor:

Basmark Corporation

     P.O. Box 40450
     Cleveland, OH 44140
     U.S.A.

Phone:

+1 216 871 8855 Fax:

+1 216 871 9011 EMail:

jgo@basmark.com URL:

http://www.basmark.com Contact:

Joseph G. O’Toole Last modified:

February 25, 1997.

10 Emulation

10.1 Emulus

Description:

Emulus is an X-Windows/Motif application which uses TCP/IP to establish a connection to an IBM mainframe host, emulating a 3270 terminal. Emulus supports mainframe graphics, including 3270 graphics datastream extension used by mainframe SAS Software. Other features include customization through dialogs, a scripting facility to record and playback host interaction, EasyPads (customized windows of buttons), and cut and paste between Emulus and other X applications. Emulus also comes with Helplus, an hypertext help viewer modeled after WinHelp. It is used to provide online help for Emulus, but can also be used to develop Help for other X applications.

Distribution media:

CD-ROM. Licensing policy:

Licensing is per CPU. OS provisions:

Linux 1.2 or higher, 8-16 MB RAM, 5 MB hard disk; Motif statically linked. Documentation:

Printed book, man pages, online help, English. Product support:

Technical Support from SAS Institute. Price range:

US$99; includes hardcopy documentation, and Technical Support for one year. Vendor:

SAS Institute Inc.

     SAS Campus Drive
     Cary, NC 27513
     U.S.A.

Phone:

+1 919 677 8000 Fax:

+1 919 677 8166 EMail:

saspjh@unx.sas.com URL:

http://www.sas.com Contact:

Phil Herold Last modified:

March 16, 1997.

10.2 Executor 2

Description:

Executor 2 reads and writes Macintosh formatted media and runs many Macintosh programs. A compatibility database is available at our web site.

Distribution media:

CD-ROM. Licensing policy:

Commercial software with a time limited crippled demo available from ftp://ftp.ardi.com or http://www.ardi.com.

OS provisions:

Linux/X/a.out, Linux/X/ELF, Linux/SVGA/a.out or Linux/SVGA/ELF. Documentation:

30 page mini-manual in CD-ROM jewel-box case. Extra features and add-ons:

Each Executor 2 CD-ROM contains Executor/DOS and Executor/Linux. You can switch between them, but are allowed to only run one at a time.

Price range:

Suggested Retail: US$249. Educational Institutions: US$149, Full-time Students: US$65. Vendor:

ARDI

     Suite 4-101
     1650 University Blvd., NE
     Albuquerque, NM 87102
     U.S.A.

Phone:

+1 505 766 9115 Fax:

+1 505 766 5153 EMail:

questions@ardi.com URL:

http://www.ardi.com Last modified:

March 19, 1997.

10.3 Wabi 2.2 for OpenLinux

Description:

Caldera licensed and ported SunSoft’s Wabi technology to enable end users to run popular Windows 3.1 applications on Linux-based system software. Channel Partners and customers can utilize Caldera Wabi 2.2 for Linux running on Caldera OpenLinux, an open source environment, to remotely manage Windows 3.1 applications at home, in the office or on the road. Channel Partners who create customized turn-key solutions based on environments like SCOR OpenServer 5 or Windows NT can now use the Wabi/COL solution to increase revenues by lowering overall cost and system requirements. The suggested retail price for Wabi 2.2 for Linux is US$199. The product requires Linux on a 386 (or higher) Intel-based processor, 16 MB RAM (24 MB recommended), VGA-quality video and 10 MB disk space.

Distribution media:

CD-ROM. Licensing policy:

One user per licensed copy. No shareware or FTP. OS provisions:

Runs on Caldera Network Desktop and OpenLinux. Documentation:

Online and printed manual. Product support:

30-days installation support via EMail and phone. Available since:

November 18, 1996. Countries with distribution:

47 countries worldwide; see our web site for local reseller. Price range:

US$299. Vendor:

Caldera, Inc.

     633 S. 550 E.
     Provo, UT 84606
     U.S.A.

Phone:

+1 800 850 7779, +1 801 377 7687 Fax:

+1 801 377 8752 EMail:

orders@caldera.com URL:

http://www.caldera.com Contact:

Keri Anderson (Sales); Nathan Hatch (Product Marketing). Last modified:

February 25, 1997.

11 Financial Software

11.1 BB Stock Pro and BB Stock Tool

Description:

BB Stock Pro and BB Stock Tool are powerful stock charting, tracking and analysis tools for investors. Features include stock charting, technical analysis, market timing, portfolio management. In addition, it has customized alerts, automatic stock split detection and management, stock screening, most active issues and big price movers, personalized watch list, Auto-Run, and builtin Easy Stock Update via Internet with either direct connection or behind firewall via proxy server.

Distribution media:

Internet (FTP and WWW). Licensing policy:

Per machine license. 30 day free trial is available at our web site. OS provisions:

Linux 2.0 ELF. Documentation:

Context sensitive online help, and User’s Guide in HTML, viewed with web browser. Product support:

Free technical support. Frequently asked questions and answers available at our web site. Available since:

Sun SPARC version available since December 1993. Linux version available since April 1995.

Countries with distribution:

USA, Canada, Sweden, Germany, Singapore, Malaysia, Taiwan. Price range:

Linux version is US$99. Vendor:

Falkor Technologies

          P.O. Box 14201
          Fremont, CA 94539
          U.S.A.

Phone:

+1 510 505 0700 EMail:

skyline@falkor.com URL:

http://www.falkor.com Contact:

Henry Chen Last modified:

March 18, 1997.

11.2 TimeClock v1.07

Description:

Employee time and attendance package for small to medium sized businesses. Includes custom report module, interface to barcode/magnetic card readers (Employee ID), and data export module.

Distribution media:

3.5” diskettes. Licensing policy:

One CPU, unlimited users. Documentation:

Printed manual and online HTML manual included with software. Product support:

Free, usually 20 minute response time.

Extra features and add-ons:

Barcode/magnetic card interface. Available since:

1992 XENIX and MS-DOS version, 1994 AIX version, October 1996 Linux version. Countries with distribution:

USA. Price range:

US$149.99. Vendor:

Quality Software Solutions, Inc.

     1322 Dodds Ave.
     Chattanooga, TN 37404
     U.S.A.

Phone:

+1 423 821 5757 Fax:

+1 423 629 0744 EMail:

dbryson@tclock.com URL:

http://www.tclock.com Contact:

Donald Bryson Last modified:

February 19, 1997.

12 Libraries

12.1 INTERACTER

Description:

INTERACTER is a portable user-interface and graphics subroutine library for Fortran software developers. INTERACTER-based programs are portable to a wide range of Fortran compilers running on Unix, VMS, Windows, DOS and OS/2. The Linux version supports f2c and g77 (ELF) and Absoft f77 (ELF). Functions include menus, text windows, forms, graphics primitives, presentation graphics, hardcopy (many formats) and operating system interface.

Distribution media:

3.5” diskettes and 4mm DAT. Licensing policy:

INTERACTER-based programs can be distributed royalty free. Linux demo program available at http://www.demon.co.uk/issltd.

OS provisions:

Linux version supports X Windows (via Xlib), graphics terminals or Linux console (colour text-mode). Requires g77, f2c or Absoft f77.

Documentation:

Two-volume English language printed manual. Product support:

Free lifetime technical support. New releases every six months. Chargeable software updates. Extra features and add-ons:

Optional INTERACTER Toolkit provides form designer, graphics file viewer, online help, hardcopy post-processor and set-up file editor.

Available since:

October, 1995. Countries with distribution:

Distributors in North America, Europe and Australia. Price range:

Linux version: US$675 (UK £450). Vendor:

Interactive Software Services Ltd.

     Westwood House
     Littleton Drive
     Huntington
     Staffs WS12 4TS
     United Kingdom

Phone:

+44 1543 503611 Fax:

+44 1543 574566 EMail:

support@issltd.demon.co.uk URL:

http://www.demon.co.uk/issltd Last modified:

February 19, 1997.

12.2 Matrix<LIB> - C++ Math Matrix Library

Description:

Matrix<LIB> is a Matlab Compatible C++ Matrix Class Library, designed for development of advanced scientific high-level C++ code. Our main design objective was to keep the C++ source code clear and similar to the equivalent Matlab code, making your shift from Matlab to C++ easier. If you know how to program Matlab, you already know how to use our Matrix<LIB> C++ Math Library. The code was carefully optimized for maximal performance, too. Matrix<LIB> is also included in the MATCOM V2 distribution. The library includes complex math, binary and unary operators, powerful indexing capabilities, signal processing, file I/O, linear algebra, string operations and graphics. Over 300 mathematical functions are included in Matrix<LIB>. MS Windows and many Unix platform are supported with a variety of C++ compilers. An evaluation version (fully functional, 30 days limited) is available for free download from our web site.

Distribution media:

Internet (WWW). Licensing policy:

From single license to site license. Contact The MathTools for details. OS provisions:

Linux 1.2, gcc 2.6.3 or later, 8 MB of RAM, 15 MB on disk. Documentation:

Users manual and library reference in various formats (MS Word, MS Write, plain text, PostScript). Product support:

A fast product support is given by EMail only. The support section in the web page give immediate access to a wealth of resources, technical support, answers to frequently asked questions, technical documents and more.

Extra features and add-ons:

MathTools Accelerator: provides significant performance gains for Matrix<LIB> programs using Linear Algebra and matrix multiplication, on a Pentium with the Visuall C++ compiler.

Available since:

November, 1996. Countries with distribution:

World wide. Price range:

US$399 (US$199 for academic institutes). Vendor:

MathTools, Ltd.

     P.O. Box 855
     Horsham, PA 19044-0855
     U.S.A.

Fax:

+1 888 MATHTOOLS, +1 215 957 1719 EMail:

info@mathtools.com URL:

http://www.mathtools.com Contact:

Robert G. Ford Last modified:

March 18, 1997.

12.3 PKWARE Data Compression Library for Linux

Description:

The library does lossless compression and consists of compression, an extraction, and an error checking routine. To maintain flexibility and keep the library’s code size at a minimum, the PKWARE Data Compression Library does not create .ZIP compatible archives.

Distribution media:

3.5” diskettes. Licensing policy:

Per machine with no runtime royalties. OS provisions:

Intel x86, 36 KB RAM for compressing and 13 KB RAM for uncompressing, 1 MB on hard disk.

Documentation:

Printed 33 page manual and separate installation notes in English. Example code provided on disk. Product support:

No-charge. Technical support via phone, fax or EMail. Extra features and add-ons:

Also compatible with FreeBSD. Available since:

September 14, 1995. Countries with distribution:

Australia, Brazil, France, Germany, India, Italy, Germany, The Netherlands, Japan, and the UK. Price range:

US$450. Educational discount available. Vendor:

PKWARE, Inc.

          9025 N. Deerwood Drive
          Brown Deer, WI 53223-2480
          U.S.A.

Phone:

+1 414 354 8699 Fax:

+1 414 354 8559 EMail:

info@pkware.com URL:

http://www.pkware.com Contact:

Bob Gorman Last modified:

March 24, 1997.

12.4 SIMLIB IG

Description:

SIMLIB IG is a C library. It enables the user to communicate with an Evans & Sutherland image generator (Liberty and ESIG Systems) using a very efficient raw Ethernet protocol. There is no need for using opcodes, since SIMLIB IG provides an API to the functionality of the image generators.

Distribution media:

4mm DAT tar format and 3.5” diskettes. Licensing policy:

Commercial. OS provisions:

Linux 2.0. Documentation:

Printed manual in English, source code examples on media. Price range:

US$2,500. Vendor:

KNIENIEDER Simulationstechnik

     Technologiezentrum Tirol
     Eduard Bodemgasse 5
     A-6020 Innsbruck
     Austria

Phone:

+43 512 390415 650 Fax:

+43 512 364000 20, +43 512 364000 30 EMail:

office@knienieder.co.at Contact:

Stephan Haidacher Last modified:

April 8, 1997.

13 Mathematics

13.1 Maple V Release 4 - The Power Edition

Description:

Maple V Release 4 is the powerful computer algebra system used by mathematicians, engineers, and scientists for teaching and research. Maple V lets you perform a wide range of symbolic and numeric computations quickly and accurately. It includes more than 2500 built-in math functions, 2D and 3D graphics, animation, and more. New features in Release 4 include outlining, styles, and hyperlinking, typeset math notation in input, output, and text, full text search of Help Database and new computational and plotting facilities.

Distribution media:

CD-ROM. Licensing policy:

License for unlimited use in time of the bought version, floating license with user based pricing. OS provisions:

Linux kernel 2.0.0 or above, X11 R6.1 or above (for running X11 GUI components), 12 MB of RAM suggested. The binaries are available in ELF format only.

Documentation:

3 books: ”Maple V - Learning Guide”, ”Maple V - Programming Guide”, ”The Maple V Handbook”. Also available in German: ”Einführung in Maple V” und ”Programmieren mit Maple V”. Documentation and help in English, a variety of books available in different languages.

Product support:

Updates/maintenance depending on license, please call the distributor or WMSI. Extra features and add-ons:

Share library via EMail and FTP. Free mailing list available (maple group@daisy.waterloo.edu). Available since:

September, 1994. Countries with distribution:

USA/Canada, Europe and Asia. Vendor:

Waterloo Maple Inc.

     450 Phillip Street
     Waterloo, ON N2L 5J2
     Canada

Phone:

+1 519 747 2373 Fax:

+1 519 747 5284 EMail:

info@maplesoft.com Vendor in Germany:

Scientific Computers GmbH

     Franzstr. 106
     D-52064 Aachen
     Germany

Phone:

+49 241 47075 0 Fax:

+49 241 44983 EMail:

maple@scientific.de Contact:

Andreas Himmeldorf Last modified:

March 26, 1997.

13.2 MATCOM and MATCOM MATH LIBRARY

Description:

MATCOM is a Matlab to C++ compiler and C++ MATRIX MATH LIBRARY. MATCOM compiles Matlab source files (M-Files) to C++ source code. The resulting code is linked with the supplied C++ Matrix class library to create standalone executables or MEX files. The C++ code and library can be integrated in products, royalty free, saving the need to translate the algorithm prototype. Compiled code runs significantly faster than the original interpreted source. Matrices of doubles, floats, ints and chars are supported, providing lower memory usage for many applications, especially signal and image processing. MS Windows and many Unix platform are supported with a variety of C++ compilers. An evaluation version (fully functional, 30 days limited) is available for free download from our web site.

Distribution media:

Internet (WWW). Licensing policy:

From single license to site license. Contact The MathTools for details. OS provisions:

Linux 1.2, gcc 2.6.3 or later, 8 MB of RAM, 15 MB on disk. Documentation:

Users manual and library reference in various formats (MS Word, MS Write, plain text, PostScript). Product support:

A fast product support is given by EMail only. The support section in the web page give immediate access to a wealth of resources, technical support, answers to frequently asked questions, technical documents and more.

Extra features and add-ons:

MathTools Accelerator: provides significant performance gains for MATCOM programs using Linear Algebra and matrix multiplication, on a Pentium with the Visuall C++ compiler.

Available since:

May, 1995. Countries with distribution:

World wide. Price range:

US$399 ($199 for academic institutes). Vendor:

MathTools, Ltd.

     P.O. Box 855
     Horsham, PA 19044-0855
     U.S.A.

Fax:

+1 888 MATHTOOLS, +1 215 957 1719 EMail:

info@mathtools.com URL:

http://www.mathtools.com Contact:

Robert G. Ford

Last modified:

March 15, 1997.

13.3 MATLAB and Simulink

Description:

MATLAB is a high-performance language for technical computing. It integrates computation, visualization, and programming in an easy-to-use environment where problems and solutions are expressed in familiar mathematical notation. MATLAB supports the development of large complex applications. Features include the M-file Performance Profiler for algorithm optimization, the integrated visual editor and debugger, and powerful interactive tools for signal processing and control system design. Simulink is a powerful, interactive software package for modeling, analyzing, and simulating dynamic systems. It provides extremely accurate simulations of nonlinear continuous-time systems and systems with multiple modes of operation. Features include support for conditionally executed subsystems; state-of-the-art solvers; usability improvements such as signal and port labeling; and new blocks.

Distribution media:

CD-ROM and Internet (FTP).

Licensing policy:

Using GlobeTrotter (FLEXlm) CPU locked or floating network options. Contact The MathWorks for details.

OS provisions:

Linux 2.0.18, 32 MB of RAM, 40 MB disk space (60 MB to include all help files and documentation).

Documentation:

Printed documentation, online help, MATLAB Reference Guide (HTML).

Product support:

The MathWorks supports a variety of Internet-based services to provide product information and services 24 hours a day. These services give immediate access to a wealth of resources, product data sheets, technical support, answers to frequently asked questions, user-contributed materials, open-forum discussions, and more.

Extra features and add-ons:

MATLAB Extensions: MATLAB Compiler, MATLAB C/C++ Math Library. Toolboxes: Fuzzy Logic, Signal Processing, Image Processing, Spline, Symbolic Math, Neural Network, Control System, Robust Control, Mu-Analysis & Synthesis, System Identification, Optimization, Statistics, Financial, Frequency Domain System Identification, Higher-Order Spectral Analysis, Communication, LMI Control, Model Predicitve Control, NAG Foundation, Neural Network, Partial Differential Equation, QFT Control Design, Symbolic Math and Extended Symbolic Math, Wavelet. Simulink Extensions: Simulink Accelerator, Real-Time Workshop. Blocksets: Fixed-Point, DSP, Nonlinear Control Design.

Available since:

June, 1995. Countries with distribution:

World wide. Price range:

Upon request. Vendor:

The MathWorks, Inc.

     24 Prime Park Way
     Natick, MA 01760
     U.S.A.

Phone:

+1 508 647 7000 x7322 Fax:

+1 508 647 7002 EMail:

efroio@mathworks.com, brian@mathworks.com URL:

http://www.mathworks.com Contact:

Enza Froio (Marketing), Brian Bourgault (Technical) Last modified:

March 27, 1997.

14 Network Servers

14.1 Aventail Internet Policy Manager

Description:

Commerical sOCKS5 server for Linux, easily extensible authentication and encryption modules. Distribution media:

/ CD-ROM and Internet (FTP). Licensing policy:

Commercial, free demo availalbe from http://www.aventail.com/download.html OS provisions:

Linux 2.0, 16 MB of RAM, amdinistration tool requires X11R5 or later. Approximately 14 MB of hard disk space required (less if you have Perl 5.003 and Perl/Tk installed).

Documentation:

English documentation available in PDF or Microsoft Word format. Product support:

90 days free technical support. Yearly maintenance is 20% of initial purchase price. Maintenance includes product upgrades.

Extra features and add-ons:

Various plugins to add functionality at run-time. Encryption/VPN plugins available domestically. Available since:

November, 1996. Countries with distribution:

World wide. VPN products avaiable only in the USA due to export restrictions. Price range:

Starting at US$1495. Vendor:

Aventail, Corp.

     117 South Main, Suite 400
     Seattle, WA 98104
     U.S.A.

Phone:

888.762.5785, 206.777.5600 Fax:

206.777.5656 EMail:

info@aventail.com URL:

http://www.aventail.com Last modified:

March 26, 1997.

14.2 Critical Angle X.500 Enabler

Description:

The Critical Angle X.500 Enabler allows an LDAP-only directory server to be integrated into an X.500 directory service, such as the PARADISE international white pages directory or an enterprise messaging directory service. It supports all X.500(88) operations, including authentication.

Distribution media:

Internet (WWW). Licensing policy:

Licensed per server. OS provisions:

Reference platform is Red Hat Linux 3 or later for Intel. Documentation:

English documentation is included. Product support:

An upgrade subscription is also available. Countries with distribution:

World wide. Price range:

A single server license is US$495. There is also (in early 1997) a special version, PARADISE X.500 Enabler, exclusively for universities. This version does not include the support for authentication or modification operations; it is intended only for searching public information. A single server license of PARADISE X.500 Enabler is currently US$95.

Vendor:

Critical Angle Inc.

     4815 W. Braker Lane \#502-385
     Austin TX 78759
     U.S.A.

EMail:

info@critical-angle.com URL:

http://www.critical-angle.com Contact:

Mark Wahl Last modified:

March 17, 1997.

14.3 DNEWS News Server

Description:

DNEWS is an advanced, full featured NNTP based news server. It is suitable for ISP’s and Corporate Intranets. DNEWS can be used as a replacement for INN or cnews and will generally result in improved performance and much lower administration. Features include: easy installation and management, flexible expire settings, flexible access restrictions, EMail confirmation of posts, XOVER extentions implemented, streaming extentions built in, dynamic sucking feeds option, news to web gateway included.

Distribution media:

Internet (WWW). Licensing policy:

Commercial, DNEWS is try before you buy software and can be downloaded from our web site for free trial.

OS provisions:

No special requirements Documentation:

A user manual is provided in HTML format. Online help included. Product support:

Supported by Netwin LTD.

Extra features and add-ons:

DNEWSWEB - A news to web gateway. Allows local and Usenet news groups to be added to ordinary HTML pages and allows users to read and post to local and Usenet news groups directly using any web browser.

Available since:

1995. Countries with distribution:

World wide. Price range:

US$485 (Free registration for schools and universities). Vendor:

Netwin LTD

     P.O Box 27574
     MT Roskill
     Auckland
     New Zealand

Fax:

+64 9 6300 689 EMail:

netwin@netwinsite.com URL:

http://netwinsite.com Contact:

Stephen Pugmire Last modified:

March 27, 1997.

15 Office Tools

15.1 The American Heritage Dictionary Deluxe

Description:

This long-awaited release is ideal for technical writers, authors, and researchers. Previously available only on PC’s, the product offers the ultimate reference tool for Linux users, with full definitions for well over 200,000 words, plus pronunciations, derivations, proper usage, idioms and sample sentences. The world reknowned Roget’s 500,000 word Thesarus is fully integrated with the AHD. Browse, word and topic search, anagrams and more. Single user, network, academic and site licenses are available for Linux and key Unix platforms.

Distribution media:

CD-ROM, tapes, Internet (FTP and WWW). Licensing policy:

Single user node locked or server license (shared). A free unclockable demo is available at our web site.

OS provisions:

Motif GUI, 8-16 MB RAM, 15 MB harddisk or run from the CD-ROM. Documentation:

Documentation as a help text file. English language. Product support:

Maintenance available. Call DUX for details. Extra features and add-ons:

Addlitional reference volumes and additional language support to be announced. Available since:

February, 1997. Countries with distribution:

Currently sold in the USA and Germany. However, sales are made internationally via unclockable FTP/WWW copies and sold with authorized major credit card.

Price range:

US$49.95 for node lock single user. Call DUX or an authorized reseller for sliding scale or site license pricing.

Vendor:

DUX Software Corporation

     425 Sherman Ave. Ste 330
     Palo Alto, CA 94306
     U.S.A.

Phone:

+1 800 543 4999, +1 415 473 1800 Fax:

+1 415 462 8723 EMail:

sales@dux.com URL:

http://www.dux.com Contact:

Bob Adams, Rob DuFrane Last modified:

March 16, 1997.

15.2 Applixware Office Suite for Linux

Description:

Fully-graphical, fully-integrated office suite for Linux. Includes word processor, spreadsheet, presentation graphics, mail and HTML authoring software. Also ships with a version of Red Hat Linux. The Developer’s Version includes all of the above and the Builder (rapid application development) program.

Distribution media:

CD-ROM. Licensing policy:

Per machine license. OS provisions:

Kernel 1.2.13 or better, XFree86 3.1.2 or better, 16 MB RAM, 150 MB hard disk. Documentation:

Printed docs approximately 750 pages, online docs approximately 3000 pages, English. Product support:

Installation support for 60 days by EMail or fax. Additional support based on number of users, etc. Contact Red Hat Software. Upgrades are offered at discounted prices to registered users. Applixware users can subscribe to applixware-list@redhat.com and archives of FAQs on web site.

Extra features and add-ons:

Add-on documentation. Available since:

July, 1996. Countries with distribution:

North America, Europe, Japan, Australia.

Price range:

US$79.95 (student version) to $495 (developer’s version). Vendor:

Red Hat Software, Inc.

     3203 Yorktown Ave., Suite 123
     Durham, NC 27713
     U.S.A.

Phone:

+1 888 REDHAT 1, +1 919 572 6500 Fax:

+1 919 572 6726 EMail:

redhat@redhat.com URL:

http://www.redhat.com Contact:

Lisa Sullivan <sulli@redhat.com> Last modified:

March 17, 1997.

15.3 D.M.S. Document Management System

Description:

The system collects documents in Tiff-g4, PostScript and JPEG and organizes them in folders with record information like names and descriptions. The system can be access using any web broswer: viewing, modify, full text retrieval and faxing are provided by a web interface. New documents are made by printing to printer spooler.

Licensing policy:

The system is installed by vendor on your Linux box with a server licence, unlimited clients, unlimited document number.

OS provisions:

Linux 2.0. Documentation:

HTML.

Extra features and add-ons:

SQL server integrationm, print capture from accounting software. Price range:

Upon request. Vendor:

Studio LEADER Pro

          Via Pietrastretta, 76
          I-38100 Trento
          Italy

Phone:

+39 461 828229 Fax:

+39 461 829862, +39 461 829877 EMail:

DMS@leader.it URL:

http://www.leader.it Contact:

Guido Brugnara Last modified:

April 11, 1997.

15.4 HotWire EasyFAX

Description:

Fully functional GUI Fax Management Program. Distribution media:

Internet (WWW). Licensing policy:

Free fully functional eval from WWW site.. OS provisions:

5.5 MB Disk, 4 MB RAM.

Documentation:

Online context-sensitive help. Product support:

Via EMail and Usenet only. Available since:

January 1, 1997. Countries with distribution:

USA/Canada (Unisource Systems, Inc.), Germany/Europe (Delix Computer, GmbH). Price range:

US$99. Vendor:

Unisource Systems, Inc.

          1409 N. Cove Blvd.
          Longwood, FL 32750
          U.S.A.

Phone:

+1 407 834 1973 Fax:

+1 407 834 8013 EMail:

sales@unisrc.com URL:

http://home.xl.ca/perfectBackup Contact:

Gary Heller Last modified:

February 14, 1997.

15.5 NExS, the Network Extensible Spreadsheet

Description:

NExS is a full-featured graphical spreadsheet package which is designed to take advantage of the advanced features of Unix and the X Window system. It includes more than 230 computational functions and 13 2D and 3D graph types for engineering, scientific and business applications. NExS is extensible through a network aware API that allows user-defined ”plug-ins” to add new features and functions in a variety of programming languages. NExS software plug-in bus includes 63 ”slots” for remote processes to connect and extend NExS’ basic functionality, allowing for the development of sophisticated networked applications.

Distribution media:

Internet (FTP and WWW).

Licensing policy:

The NExS Personal Edition license is for a single user and single platform type. The owner may use the license on multiple machines. The Professional Edition license is a floating network license which allows a pool of licenses to be used on any supported platform in your network. Without a license, NExS runs in ”demo mode” with limited functionality.

OS provisions:

NExS runs on Linux kernels which support ELF binaries. There are two flavors for Linux, one which requires Motif 2.0 shared libraries, and one which includes Motif 2.0 statically linked.

Documentation:

The NExS User Guide is available in PostScript (letter format, 431 pages, or A4 format, 403 pages), and in HTML. API and tclNExS documentation are available separate from the User Guide, also in PostScript and HTML format. Printed documentation is also available. All documentation is in English.

Product support:

EMail support is standard at no cost. Phone-in support is available for a charge. Bug fix and minor version releases (e.g., 1.2.x -¿ 1.3.0) are no cost. Major releases are substantially discounted for existing licensees.

Extra features and add-ons:

NExS plug-ins and user-contributed code are available at http://www.xess.com/plugins.html.

Available since:

1994.

Countries with distribution:

World wide.

Price range:

• US$149 ($99 academic) for Personal Edition

• US$249 ($125 academic) for Professional Edition

Vendor:

X Engineering Software Systems (XESS) Corp.

     P.O. Box 33091
     Raleigh, NC 27636
     U.S.A.

Phone:

+1 800 961 7840, +1 919 387 0076 Fax:

+1 919 387 1302 EMail:

info@xess.com URL:

http://www.xess.com FTP:

ftp://ftp.vnet.net/pub/users/xess/NExS Contact:

Dave Van den Bout <devb@xess.com> Last modified:

March 15, 1997.

16 Text Processing

16.1 Edith Pro for X11

Description:

Edith Pro is a powerful and easy-to-use text editing package, for editing plain text such as EMail messages, HTML, LaTeX and programming languages. It is aimed at people who want a flexible editing tool but who are not prepared to spend much time into learning and configuring their application. It includes a fast text-only web browser with access to the system manual and info.

Distribution media:

Internet (WWW).

Licensing policy:

You are only allowed to download Edith from the ZFC web pages for personal use. Without a licence, Edith will regularly issue messages stressing that the copy used is not licenced. Personal use without a licence is allowed, but organizations are required to buy a licence.

OS provisions:

ELF binary. 1MB is sufficient for editing normally sized text. Documentation:

Online HTML help covers the printed manual completely. The printed manual contains illustrations not in the online help.

Product support:

Free updates can be downloaded from the ZFC web pages. Licence holders will be notified of updates. Licence holders can contact ZFC by EMail for immediate support.

Available since:

January 1, 1997. Countries with distribution:

World wide directly from ZFC. Price range:

NLG 100 single host, NLG 500 maximum. 50% discount for educational organizations and students. Vendor:

ZFC

     P.O. Box 15813
     1001 NH Amsterdam
     The Netherlands

Phone:

+31 20 4 208 248 EMail:

zfc@zfc.nl URL:

http://www.nl.net/˜ zfc/index.html Contact:

Annius V. Groenink Last modified:

March 17, 1997.

17 System Administration Tools

17.1 PerfectBACKUP+

Description:

PerfectBACKUP+ is the Unix/Linux world’s premier Backup/Restore utility. It is the worlds fastest backup/restore for Unix and has been since 1985. Previously sold in the US as FASTBACK PLUS for Unix, the counterpart to the DOS FASTBACK PLUS, it now has character and Motif interfaces, networking support, compression, the best verification and recovery, scheduling and is compatible with both tar and cpio. Backs up all files, devices, partitions etc. and can backup Windows, Netware and WindowsNT drives too.

Distribution media:

3.5” diskettes, Internet (WWW), 4/8mm DAT and soon CD-ROM. Licensing policy:

Free evaluation copy from http://home.xl.ca/perfectBackup. Free online manual from same the same address. Price per CPU, unlimited users.

OS provisions:

Kernels 1.2.x and 2.0.x, X11R6, Motif 1.2, 3 MB of RAM, 3 MB disk. Documentation:

Printed (188 pages), PostScript online, and online help. English language documentation, but help by phone and EMail also avilable in Spanish.

Product support:

Unlimited free EMail support and phone support. Extra features and add-ons:

Comes with free PC program to read tar/cpio diskettes onto PC. AutoChanger add-on with tape recycling for high-end shops.

Available since:

1996. Countries with distribution:

USA, Canada, Germay, UK, Switzerland, Singapore, Malaysia, Germany. Price range:

UD$79 to US$499. Linux off the net (no media) is US$79 and HP/UX is US$499. Always full product.

Vendor:

Unisource Systems, Inc.

     1409 North Cove Blvd.
     Longwood, FL 32750
     U.S.A.

Phone:

+1 407 834 1973 Fax:

+1 407 834 1973 EMail:

sales@unisrc.com URL:

http://home/xl.ca/perfectBackup Contact:

Gary Heller Last modified:

March 15, 1997.

17.2 Venus

Description:

Venus is an distributed administration tool for any network built of Unix workstations from different hardware vendors. A Venus network presents itselfs to the user as a homogeneous computer system and guarantees transparent access to any network resource. It provides software tools for major cluster management tasks: configuration management, software distribution, user and filesystem administration.

Licensing policy:

Commercial. OS provisions:

Currently build on top of the S.u.S.E distribution. Documentation:

Printed doumentation, online help. Product support:

Hotline, EMail, courses. Available since:

Countries with distribution:

World wide. Price range:

Upon request. Special prices for educational institutions. Vendor:

science + computing GmbH

     Hagellocher Weg 71
     D-72070 T\"{u}bingen
     Germany

Phone:

+49 7071 9457 0 Fax:

+49 7071 9457 27 EMail:

info@science-computing.de URL:

http://www.science-computing.de Contact:

Olaf Flebbe Last modified:

February 25, 1997.

18 X Windows Related Products

18.1 AcceleratedX Server

Description:

Commercial grade X display server with workstation-class features like multiple visuals, overlays, and gamma correction for FreeBSD and Linux. The AcceleratedX server drives the most popular graphics hardware, including the latest cards from ATI, Diamond, Matrox, NeoMagic, Number 9, and S3. In all, more than 450 different cards are supported.

Distribution media:

3.5” diskettes.

Licensing policy:

Per CPU. OS provisions:

No known problems with any Linux distribution. Documentation:

A 100 page perfect bound manual is included. Product support:

Installation support in included. Available since:

1994. Countries with distribution:

World wide. Price range:

Contact your local reseller, or Xi Graphics. Vendor:

Xi Graphics

          1801 Broadway, Suite 1710
          Denver, CO 80202
          U.S.A.

Phone:

+1 303 298 7478 Fax:

+1 303 298 1406 EMail:

sales@xig.com URL:

http://www.xig.com Contact:

Richard Van Dyke Last modified:

March 14, 1997.

18.2 BXwidgets

Description:

BXwidgets is a supplementary widget set to develop especially commercial Motif applications. It includes special widgets for tabulars, masks, formatted input field and the full support for a complete help system. All has Motif look and feel and is platform independent.

Distribution media:

3.5” diskettes. Licensing policy:

Source code licence. OS provisions:

Any kernel, X11R5 or X11R6, Motif 1.1 or Motif 1.2. Documentation:

About 250 pages of documenation with examples. Several demo programs. Product support:

Support and update service is available (DM 1200; approximately US$750) Available since:

1992. Countries with distribution:

Germany, UK and USA. Price range:

DM 1600 (approximately US$1000) for a full source code licence. Vendor:

BREDEX GmbH

     Fallersleber-Tor-Wall 23
     D-38100 Braunschweig
     Germany

Phone:

+49 531 24 33 0 0 Fax:

+49 531 24 33 0 99 EMail:

info@bredex.de URL:

http://www.bredex.de Last modified:

March 17, 1997.

18.3 BXwidgets/DB

Description:

BXwidgets/DB is a supplementary widget set to access relational databases. Data can be automatically queried form the database, displayed in various forms, changed and stored in the database. All applications built with BXwidgets/DB are database independent. The database access is performed by transaction-servers. Transaction-Server and database may reside on other hotsts in the network.

Distribution media:

CD-ROM. Licensing policy:

Per OS license, no run-time restrictions. Evaluation license available. OS provisions:

X11R5 and Motif 1.2 must be installed. Documentation:

Printed manual in English. Product support:

Support and update service is available. Available since:

1996. Price range:

DM 4900 (approximately US$2900). Vendor:

BREDEX GmbH

     Fallersleber-Tor-Wall 23
     D-38100 Braunschweig
     Germany

Phone:

+49 531 24 33 0 0 Fax:

+49 531 24 33 0 99 EMail:

info@bredex.de URL:

http://www.bredex.de Last modified:

March 17, 1997.

18.4 AcceleratedX CDE Business Desktop

Description:

A complete networked operating system that enables administrators to integrate industry-standard PCs into existing Unix networks with better internationalization and the same functionality as more expensive workstations.

Distribution media:

CD-ROM. Licensing policy:

Per CPU. OS provisions:

No known problems with any Linux distribution. Documentation:

A 100 page perfect bound manual the AcceleratedX server and a 300 page perfect bound manual for CDE is included.

Product support:

Installation support in included. Available since:

1997. Countries with distribution:

World wide. Price range:

Contact your local reseller, or Xi Graphics. Vendor:

Xi Graphics

          1801 Broadway, Suite 1710
          Denver, CO 80202
          U.S.A.

Phone:

+1 303 298 7478 Fax:

+1 303 298 1406 EMail:

sales@xig.com URL:

http://www.xig.com Contact:

Richard Van Dyke Last modified:

March 14, 1997.

18.5 AcceleratedX OSF-Certified Motif

Description:

A complete OSF-Certified Motif runtime and development environment. Distribution media:

CD-ROM. Licensing policy:

Per CPU. OS provisions:

No known problems with any Linux distribution. Documentation:

Release note that include references to well-written commerical reference manuals. Product support:

Installation support in included. Available since:

Countries with distribution:

World wide. Price range:

Contact your local reseller, or Xi Graphics. Vendor:

Xi Graphics

          1801 Broadway, Suite 1710
          Denver, CO 80202
          U.S.A.

Phone:

+1 303 298 7478 Fax:

+1 303 298 1406 EMail:

sales@xig.com URL:

http://www.xig.com Contact:

Richard Van Dyke Last modified:

March 14, 1997.

18.6 AcceleratedX SGI-Certified OpenGL

Description:

A complete SGI-Certified OpenGL server extension. Distribution media:

3.5” diskettes. Licensing policy:

Per CPU. OS provisions:

No known problems with any Linux distribution.

Documentation:

Release note that include references to well-written commerical reference manuals. Product support:

Installation support in included. Available since:

1995. Countries with distribution:

World wide. Price range:

Contact your local reseller, or Xi Graphics. Vendor:

Xi Graphics

          1801 Broadway, Suite 1710
          Denver, CO 80202
          U.S.A.

Phone:

+1 303 298 7478 Fax:

+1 303 298 1406 EMail:

sales@xig.com URL:

http://www.xig.com Contact:

Richard Van Dyke Last modified:

March 14, 1997.

19 Other Software

19.1 Clustor

Description:

Clustor is a program for managing large computational tasks. Clustor greatly simplifies a common computationally intensive activity - running the same program code numerous times with different inputs. Clustor provides increased performance by distributing jobs over a network of computers and improved task management through a friendly user interface. Targeted at users who run computationally demanding tasks, Clustor supports all phases of running such a task: task preparation, job generation and job execution. With Clustor, users who do not have extensive knowledge of programming parallel applications, such as scientists, engineers, researchers, are able to utilize the power of networked computers.

Distribution media:

Internet (WWW). Optional CD-ROM. Licensing policy:

A free demo licenses with limited functionality, 30 days evaluation licenses with complete functionality from our web site.

OS provisions:

Linux 1.2 or higher. Documentation:

Web site, optional printed documentation. Product support:

EMail support. Extra features and add-ons:

Features for load monitoring and resource sharing, network installation program. Available since:

February, 1997. Countries with distribution:

World wide Price range:

Starting at US$998. Vendor:

Active Tools Inc.

     246 First St, Suite 310
     San Francisco, CA 94105
     U.S.A.

Phone:

+1 415 882 7062 Fax:

+1 415 680 2369 EMail:

info@activetools.com URL:

http://www.activetools.com Contact:

Rok Sosic, Sergij Foski Last modified:

March 17, 1997.

19.2 FootPrints

Description:

The first web-based helpdesk software system. Tracks problems/solutions, trouble tickets – whatever information you need to manage – and makes it available via the World Wide Web to your customers or users using their web browser.

Distribution media:

3.5” diskettes, tape and Internet (FTP and WWW). Licensing policy:

Free eval available from http://www.unipress.com/cgi-bin/free evals or ftp://ftp.unipress.com/pub/free evals. Pricing is on a per-user basis.

OS provisions:

Runs with any Linux running a web server. Program size is 100K, 3MB diskspace. Documentation:

Online manual/help. Product support:

Free maintenance through May 31, 1997. 15% thereafter for maintenance and updates. Support is free.

Available since:

January 1, 1997. Countries with distribution:

World wide. Price range:

US$1995 per starter pack (server software plus 3 user licenses). Vendor:

UniPress Software, Inc.

     2025 Lincoln Highway
     Edison, NJ 08817
     U.S.A.

Phone:

+1 800 222 0550, +1 908 287 2100 Fax:

+1 908 287 4929 EMail:

info@unipress.com URL:

http://www.unipress.com Contact:

Sue Glassberg <sue@unipress.com> Last modified:

March 5, 1997.

19.3 Aladdin Ghostscript

Description:

PostScript, PDF, PCL5e (2Q97), PCL5c (2Q97), and PCL XL interpreter. Provides X Windows previewing, conversion to half a dozen raster file formats, and output to dozens of printers. Runs in all OS environments, not just Linux, and on any processor (32- or 64-bit) with a C compiler. Can drive black-and-white, RGB, and CMYK devices, both bilevel and continuous-tone. For more information, see http://www.cs.wisc.edu/˜ ghost/index.html. Note that Aladdin Ghostscript is an OEM product only: end-user licensing is not available (and not necessary, since the product is free for end users), nor is end-user support.

Distribution media:

3.5” diskettes. Licensing policy:

Normally per-CPU. License normally includes source code, upgrades, and support. OS provisions:

X11R5 or later is required for X Windows previewing. OS-specific modules are available for Unix (including Linux), VMS, MS Windows (including Windows NT), OS/2, and Macintosh. Porting to other OSs is simple: all OS dependencies are segregated in a single file of typically less than 100 lines of code.

Documentation:

Online documentation only, in English. Currently in plain ASCII text; texinfo and HTML versions are in preparation.

Product support:

Normally included in license; by EMail, phone, and fax. Extra features and add-ons:

Each input language is priced separately, with a discount for multiple languages. Available since:

1988 as free software, 1991 commercially. Countries with distribution:

World wide. Price range:

Negotiated for each customer. Vendor:

Artifex Software Inc.

     454 Las Gallinas Ave., Suite 108
     San Rafael, CA 94903
     U.S.A.

Phone:

+1 415 492 9861 Fax:

+1 415 492 9862 EMail:

info@arsoft.com URL:

http://www.cs.wisc.edu/ ghost/index.html Last modified:

February 21, 1997.

19.4 LjetMgr

Description:

Printers from Hewlett-Packard (both laser and ink variants) have an extensive set of configurable options. These options are normally set on the front panel of the printer. The most notable options are:

• Ccono-mode to save toner

• Powersave mode

• Page format selection

• Font/typeface and language selection

With LjetManager you can directly modify these settings on your screen. It is especially usefull for printers that lack front panel configurability (like the LaserJet 5L) or if you have more than one printer at your disposal. LjetManager sports a graphical user interface, is fully localizable and comes with documentation and help pages in HTML format.

Distribution media:

3.5” diskettes, Internet (EMail and WWW). Licensing policy:

You can download a demo version (in fact it’s the previous version of LjetMgr) from the following locations and their mirror sites. Just look for LjetMgr2.4.1.README and the acompanying files.

• http://www.neosoft.com/tcl/ftparchive/sorted/print

• ftp://ftp.neosoft.com/pub/tcl/sorted/print

• ftp://sunsite.unc.edu/pub/Linux/system/Printing

Note that the commercial version contains a lot of enhancements and is faster. OS provisions:

Needs a printer that supports PJL. Available for Linux on Intel and Alpha. Documentation:

Complete context sensitive Help and documentation in HTML format. A PostScript version is available on request.

Product support:

Prices include installation support and one year of free upgrades. Additional support available on request.

Available since:

May, 1996. Countries with distribution:

World wide. Price range:

A single license of LjetMgr costs US$65. Discount for educational institutions and students is 10%. This price includes installation support and one year of free upgrades. Contact us for discounts on network licenses.

Vendor:

softWorks, Richard Schwaninger

     Theodor-K\"{o}rnerstr. 173
     A-8010 Graz
     Austria

Phone:

+43 316 686590 Fax:

+43 316 686590 EMail:

risc@ping.at URL:

http://members.ping.at/risc Contact:

Richard Schwaninger Last modified:

March 26, 1997.

20 Free Software for Commercial Hardware

This section covers free Linux software (e.g. drivers) for specialized hardware.

20.1 Stallion Technologies Multiport Serial Boards

Description:

Stallion provides Linux support for its EasyIO and EasyConnection range of multiport cards. EasyIO is available with either 4 asynchronous serial ports with RJ45 connectors, or 8 asynchronous serial ports with RJ45 or DB25 connectors. EasyIO supports data rates of up to 145 Kbits/sec or 460 Kbits/sec per port (depending on the model) and full modem signaling. Drivers are available for most Intel-based operating systems, including Linux. EasyIO is available for the ISA bus. EasyConnection is an expandable multiport solution that provides from 8 to 64 asynchronous serial ports in 8 or 16 port increments. Both RJ45 and DB25 connectors are available. Features include full surge suppression, full modem signaling, and data rates of up to 145 Kbits/sec or 460 Kbits/sec per port (depending on the panel version). Drivers are available for most Intel-based operating systems, including Linux. EasyConnection is available for the ISA, EISA and PCI bus.

Distribution media:

Internet (FTP).

Licensing policy:

The Linux driver is available at no charge and is copyrighted under the GNU General Public License. Source code is provided.

OS provisions:

The most current driver supports all 2.0.x kernel versions. The drivers are updated regularly to keep pace with kernel changes.

Documentation:

Hardware reference documentation provided with the boards. The Linux driver package contains a README file with driver installation and usage information.

Product support:

Stallion Technologies now fully supports the Stallion board Linux drivers. Software updates are available from the Internet. Stallion also provides technical support for the hardware. Contact Stallion support for assistance.

Available since:

October, 1994.

Countries with distribution:

World wide distribution. Available in most countries.

Price range:

Prices for the serial boards vary from country to country. Contact your nearest Stallion Technologies office for local availability and pricing. The driver package contains a file with contact information for all Stallion Technologies offices.

Vendor:

Stallion Technologies Pty Ltd.

          2880 Research Park Drive, Suite 160
          Soquel, CA 95073
          U.S.A.

Phone:

+1 800 347 7979 Fax:

+1 408 477 0444 EMail:

info@stallion.com URL:

http://www.stallion.com FTP:

ftp://ftp.stallion.com Last modified:

February 20, 1997.